### v3

# Formal extension of noncommutative tensor-triangular support varieties

Abstract.
Given a support variety theory defined on the compact part of a monoidal triangulated category, we define an extension to the non-compact part following the blueprint of Benson–Carlson–Rickard, Benson–Iyengar–Krause, Balmer–Favi, and Stevenson. We generalize important aspects of the theory of extended support varieties to the noncommutative case, and give characterizations of when an extended support theory detects the zero object, under certain assumptions. In particular, we show that when the original support variety theory is based on a Noetherian topological space, detects the zero object, satisfies a generalized tensor product property, and comes equipped with a comparison map, then the extended support variety also detects the zero object. In the case of stable categories of finite tensor categories, this gives conditions under which the central cohomological support admits an extension that detects the zero object, confirming part of a recent conjecture made by the second author together with Nakano and Yakimov.
Key words and phrases:
noncommutative Balmer spectrum, thick ideal classification, support variety
1991 Mathematics Subject Classification:
18G65, 18G80, 18M05,
1. Introduction
Since the early 80s, support varieties have been an important tool in representation theory, when cohomological support and rank support were introduced to provide geometric techniques for studying categories of algebraic origin [AS82, Car83, Alp87]; analogous versions of support in commutative algebra and algebraic geometry were being developed around the same period [Hop87, Avr89, Nee92a]. A support theory here and throughout means a map which assigns to each object of some category a closed subset of some topological space, in a way which is compatible with the structure of the category. By the early 2000s, rank varieties were generalized to finite group schemes via the $\pi$-points of Friedlander–Pevtsova [FP07]. With the advent of tensor-triangular geometry, a new support theory (the Balmer support, universal support, or tensor-triangular support) based on the Balmer spectrum entered the literature [Bal05]. It has found widespread applications in algebraic topology, commutative algebra, algebraic geometry, modular representation theory, equivariant stable homotopy theory, and motivic homotopy theory (see [Bal20] and the references therein). Noncommutative Balmer spectra were developed in [NVY22a] to study monoidal triangulated categories without braidings, which arise naturally from various settings such as stable categories of finite tensor categories (e.g.  the representation theory of finite-dimensional Hopf algebras), group actions on tensor-triangulated categories via the crossed product (see e.g. [HV23]), and derived categories of bimodules over a noncommutative ring (see [BKSS20, Section 1]).
To greater and lesser extents, these theories of support varieties were developed to understand “sufficiently small” tensor-triangulated categories, that is, categories of compact objects. From the beginning (e.g.  when cohomological and rank support were used by Benson–Carlson–Rickard in the 90s to classify the thick ideals of the stable category of finite-dimensional modules for a finite group in characteristic $p$ [BCR97]), it has been clear that even to understand these “small” categories, it is sometimes necessary to extend a support theory to arbitrary “big” objects. For example, in the finite group case, this means extending to infinite-dimensional representations. Fundamentally, this is because Brown representability guarantees that functors preserving set-indexed coproducts have right adjoints [BIK12, Corollary 2.13]; but this construction only works when the category is compactly-generated, and in particular requires the existence of arbitrary set-indexed coproducts.
The first extended support varieties, designed to study infinitely-generated representations of a finite group over a field of positive characteristic, were introduced by Benson–Carlson–Rickard in [BCR95, BCR96], which used the Rickard idempotent functors developed in [Ric97]. This extended support variety theory was an extension of the rank / cohomological support. This approach was generalized by Benson–Iyengar–Krause to any triangulated category with an action by a graded-commutative Noetherian ring, which plays the role of the cohomology ring [BIK08, BIK12]. Shortly after the introduction of the Balmer support for symmetric tensor-triangulated categories, an analogous extension, also using versions of Rickard idempotents, was produced by Balmer–Favi [BF11]. This version of support was further developed and extended to tensor-triangulated module categories by Stevenson [Ste13, Ste18]. Beyond its use in classifying thick ideals, extended supports have also been used in the wide-ranging program of stratification, i.e.  in classifying localizing subcategories [BIK11, BIK11, Ste13, Ste14b, Ste17, BIKP17, BIKP18, BHS23a, BHS23b].
We are interested in developing the analogous theory in the context of stable categories of finite tensor categories. Finite tensor categories are abelian monoidal categories satisfying finiteness properties; they have played key roles in representation theory, mathematical physics, and topology (see [EGNO15]). Their stable categories are (possibly noncommutative) monoidal triangulated categories, and there is significant attention currently being given to support varieties, and in particular in producing a concrete support datum that realizes their Balmer support, see [BPW21, NVY22a, NVY22b, NP23a, NP23b, NVY24]. While a version of extended support for the noncommutative Balmer support was given in [NVY24, Appendix B], there has not yet been an attempt to systematically develop extended support based on Rickard idempotents for supports other than the Balmer one. Developing such a theory would be advantageous, by a theorem of the second author with Nakano and Yakimov, which states that if $\sigma$ is a support map on a monoidal triangulated category $\mathbf{K}^{c}$ with value in a Zariski space $X$ which satisfies faithfulness, realization, and a tensor product property, and admits an extension which also satisfies faithfulness and a tensor product property, then $\operatorname{Spc}\mathbf{K}^{c}\cong X$ [NVY22a, Theorem 1.4.1]. We give a precise formulation of this theorem in Theorem 2.9, after having recalled the necessary technical background in Section 2.
There is a candidate for a support which realizes the Balmer support for the stable category of a finite tensor category, namely the central cohomological support $\operatorname{supp}^{\rm C}$ introduced by the second author with Nakano and Yakimov [NVY24, Section 1.4]. This support datum is defined similarly to classical cohomological support, but is based on a particular subalgebra of the cohomology ring, rather than the whole cohomology ring. It is conjectured in [NVY24, Conjecture E] that the central cohomological support satisfies the conditions of Theorem 2.9, which would imply that $\operatorname{Spc}\underline{\mathbf{C}}$ is homeomorphic to the projective spectrum of the ring $C^{\bullet}$.
In particular, [NVY24, Conjecture E] conjectures that the central cohomological support admits a faithful extension which satisfies the generalized tensor product property. The goal of this paper is to develop a general extension for an arbitrary support datum in the noncommutative situation based on Rickard idempotent functors, and determine the precise extent to which the existence of a faithful extension satisfying the tensor product property in the situation of Theorem 2.9 follows from the other assumptions of that theorem. We believe that this represents a foundational question and is a necessary step in developing the theory of noncommutative support varieties. We first find the following:
Theorem 1.1 (Theorem 5.10, Corollary 6.6, Theorem 7.6).
Let $\mathbf{K}$ be a rigidly-compactly-generated monoidal triangulated category with compact part $\mathbf{K}^{c}$ which is either $\operatorname{Spc}$-Noetherian or has a thick generator. Let $(X,\sigma)$ be a support datum (see Section 2) defined on $\mathbf{K}^{c}$, and let $\sigma$ be faithful (that is, $\sigma(A)=\varnothing$ if and only if $A\cong 0$).
  1. (1)
    If $\sigma$ is tensorial ($\bigcup_{B\in\mathbf{K}^{c}}\sigma(A\otimes B\otimes C)=\sigma(A)\cap\sigma(B)$ for all $A$ and $B\in\mathbf{K}^{c}$) and $X$ is a Zariski space (Noetherian and sober), then $\sigma$ admits an extension $\widetilde{\sigma}$ to $\mathbf{K}$. Suppose $\sigma$ is comparative, that is, the universal map $\eta:X\to\operatorname{Spc}\mathbf{K}^{c}$ admits a left inverse $\rho$ such that $\eta\rho(\mathbf{P})\subseteq\mathbf{P}$ for all $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$. Then the extension $\widetilde{\sigma}$ is faithful. When $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian, the opposite implication holds as well, and the following conditions are equivalent:
    1. (a)
      $\sigma$ is comparative;
    2. (b)
      every closed subset $V$ of $X$ has the form $\sigma(A)$ for some $A\in\mathbf{K}^{c}$, and for each $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$, the closed set $\eta^{-1}(\overline{\{\mathbf{P}\}})$ in $X$ has a unique generic point;
    3. (c)
      the extension $\widetilde{\sigma}$ of $\sigma$ is faithful.
  2. (2)
    If $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian and if there exists a closed surjective map $\rho:\operatorname{Spc}\mathbf{K}^{c}\to X$ such that $\sigma(A)=\rho(\operatorname{supp}^{\rm B}(A))$ for all $A\in\mathbf{K}^{c}$ (where $\operatorname{supp}^{\rm B}$ denotes the Balmer support) then $\sigma$ admits a faithful extension $\widetilde{\sigma}$ to $\mathbf{K}$.
In the case of central cohomological support for finite tensor categories, we are able to use Theorem 1.1 in combination with results from [NVY24] to prove:
Theorem 1.2 (Theorem 8.6).
Let $\mathbf{C}$ be a finite tensor category with stable category $\underline{\mathbf{C}}$ and categorical center $C^{\bullet}$ such that $\operatorname{Ext}^{\bullet}_{\mathbf{C}}(A,A)$ is finitely-generated as a $C^{\bullet}$-module, for each $A\in\mathbf{C}$. Then the central cohomological support $\operatorname{supp}^{\rm C}$ admits a faithful extension to $\underline{\operatorname{Ind}(\mathbf{C})}$ if either of the following conditions holds:
  1. (1)
    $\operatorname{Proj}C^{\bullet}$ is Noetherian and central cohomological support is tensorial;
  2. (2)
    $\operatorname{Spc}\underline{\mathbf{C}}$ is Noetherian and $C^{\bullet}$ is finitely-generated.
This shows that one of the facets of [NVY24, Conjecture E] follows from the other conjectured results.
The paper is organized as follows: in Section 2 we review background on monoidal triangular geometry and support varieties. In Section 3, we define extended support via Rickard idempotents in the noncommutative case, based on the classical commutative constructions. In Section 4, we prove that the extension of the Balmer support satisfies a generalized tensor prodouct property, and along the way give some useful descriptions of the extended Balmer supports of the Rickard idempotents themselves. In Section 5, we prove that any extension of a support equipped with a comparison map (that is, a map of topological spaces with compatibility properties that allow a “comparison” between the support in question and the Balmer support) is faithful, as long as either the Balmer spectrum of $\mathbf{K}^{c}$ is Noetherian or $\mathbf{K}^{c}$ is has a thick generator. In Section 6, we show that if the Balmer spectrum is Noetherian, then any faithful and realizing support which satisfies a generalized tensor product property and admits a faithful extension also admits a comparison map. In Section 7, we prove that if a support map is induced by a surjective map from a Noetherian Balmer spectrum, then the extension of that support is faithful. Lastly, in Section 8, we turn our attention to the specific case of stable categories of finite tensor categories, and prove Theorem 1.2, that is, that extended central cohomological support is faithful under certain assumptions. We then apply these results to answer a question of Pevtsova–Witherspoon [PW15], which recovers using different techniques a theorem of Negron–Pevtsova [NP23a].
Remark 1.3 (Note on further applications).
Beyond the applications to finite tensor categories presented in Section 8, the results of this paper are applied in joint work of the second author with D.K.  Nakano and M.T.  Yakimov [NVY25] to study stratification in monoidal triangular geometry and prove a version of Balmer’s Nerves of Steel Conjecture for crossed product categories. Future applications will also include work currently in preparation joint between the second author with Ø.  Solberg and S.  Witherspoon to study Balmer spectra of some monoidal triangulated categories related to bimodules for self-injective algebras.
Acknowledgements
We thank Greg Stevenson for helpful suggestions which improved the paper. The first author was partially supported by NSF graduate student fellowship DGE-2140743. The second author was partially supported by NSF postdoctoral fellowship DMS-2103272 and by an AMS-Simons Travel Grant.
2. Monoidal triangular geometry and support theories
In this section, we review the basic background of monoidal triangular geometry. For additional details on classical (commutative) tensor-triangular geometry see [Bal05]. For details about the noncommutative generalization, see [NVY22a].
Let $\mathbf{K}$ be a triangulated category, with shift functor $\Sigma$. For background on triangulated categories, see [Nee01]. Recall that a triangulated subcategory is one closed under cones and $\Sigma^{\pm}$; a thick subcategory is a full triangulated subcategory closed under direct summands; a localizing subcategory is a full thick subcategory closed under arbitrary set-indexed coproducts (if such coproducts exist). We briefly remark that thickness follows automatically from the other properties of localizing subcategories as long as $\mathbf{K}$ is closed under set-indexed coproducts via the Eilenberg swindle, but the thickness hypothesis is important if such coproducts don’t exist.
Throughout, we assume that $\mathbf{K}$ is a rigidly-compactly-generated monoidal triangulated category. To spell this out:
  1. (1)
    $\mathbf{K}$ is triangulated, and contains all set-indexed coproducts;
  2. (2)
    $\mathbf{K}$ is generated, as a localizing subcategory, by $\mathbf{K}^{c}$, the subcategory of $\mathbf{K}$ consisting of compact objects (that is, $\mathbf{K}$ is the smallest localizing subcategory of $\mathbf{K}$ which contains $\mathbf{K}^{c}$);
  3. (3)
    $\mathbf{K}$ is equipped with a biexact tensor product tensor-product-$\otimes$-- ⊗ - with tensor unit $\mathbf{1}$ satisfying the usual associativity and unit axioms [EGNO15, Definition 2.1.1], and the tensor product commutes with arbitrary set-indexed coproducts;
  4. (4)
    $\mathbf{1}\in\mathbf{K}^{c}$, and if $A$ and $B$ are in $\mathbf{K}^{c}$, then so is $A\otimes B$ (in particular, $\mathbf{K}^{c}$ is a tensor subcategory of $\mathbf{K}$);
  5. (5)
    $\mathbf{K}^{c}$ is rigid, as a monoidal category (that is, every compact object has a left and right dual, [EGNO15, Definition 2.10.1, 2.10.2]).
Note that we do not assume the tensor product is symmetric or braided. Our primary motivating example, which we explore in more detail in Section 8, is the case of stable categories of finite tensor categories.
Definition 2.1.
We call a monoidal triangulated category big if it satisfies (1)-(5) above, that is, if it is a rigidly-compactly-generated monoidal triangulated category. We call a monoidal triangulated category small if it is equivalent to $\mathbf{K}^{c}$ for some big monoidal triangulated category $\mathbf{K}$. We call $\mathbf{K}^{c}$ finitely-generated if there exists an object $G$ in $\mathbf{K}^{c}$ such that the smallest thick subcategory containing $G$ is $\mathbf{K}^{c}$.
Recall that a thick ideal of $\mathbf{K}^{c}$ is a thick subcategory closed under tensor products with arbitrary objects of $\mathbf{K}^{c}$. Given a collection of objects $\mathbf{S}$ of $\mathbf{K}^{c}$, we will denote $\langle\mathbf{S}\rangle$ the smallest thick ideal containing $\mathbf{S}$. Likewise, we denote by $\operatorname{Loc}(\mathbf{S})$ and $\operatorname{Loc}^{\otimes}(\mathbf{S})$ the smallest localizing subcategory and smallest localizing ideal of $\mathbf{K}$ generated by any collection of objects $\mathbf{S}$ in $\mathbf{K}$. A proper thick ideal $\mathbf{P}$ is called prime if
$\mathbf{I}\otimes\mathbf{J}\subseteq\mathbf{P}\Rightarrow\mathbf{I}\text{ or }\mathbf{J}\subseteq\mathbf{P}$
for all thick ideals $\mathbf{I}$ and $\mathbf{J}$. The category $\mathbf{K}^{c}$ is called completely prime if every prime ideal is completely prime, that is, for a prime ideal $\mathbf{P}$, we have $A\otimes B$ in $\mathbf{P}$ implies $A$ or $B$ in $\mathbf{P}$, for any objects $A$ and $B$ in $\mathbf{K}^{c}$. The Balmer spectrum $\operatorname{Spc}\mathbf{K}^{c}$ is defined as the topological space consisting of prime thick ideals of $\mathbf{K}^{c}$, with closed sets
$\operatorname{supp}^{\rm B}(\mathbf{S}):=\{\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}:\mathbf{P}\cap\mathbf{S}=\varnothing\},$
for any collection $\mathbf{S}$ of objects of $\mathbf{K}^{c}$. Appropriately, we call $\mathbf{K}^{c}$ $\operatorname{Spc}$-Noetherian if $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian. The Balmer spectrum is useful conceptually because under mild conditions, it gives a topological classification of the thick ideals of $\mathbf{K}^{c}$ by Thomason-closed subsets of $\operatorname{Spc}\mathbf{K}^{c}$. In general, for a topological space $X$, $\operatorname{Thom}(X)$ will denote the set of Thomason-closed subsets of $X$ (recall that a set is Thomason-closed if it is a union of closed sets with quasicompact complement). The classification theorem is given by the following:
Theorem 2.2.
When $\mathbf{K}^{c}$ is commutative, completely prime, $\operatorname{Spc}$-Noetherian, or finitely-generated, there are inverse bijections between the thick ideals of $\mathbf{K}^{c}$ and the Thomason-closed subsets of $\operatorname{Spc}\mathbf{K}^{c}$ given by
${\mathbf{I}}$${\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I}):=\bigcup_{A\in\mathbf{I}}\operatorname{supp}^{\rm B}(A)}$${\operatorname{Ideals}(\mathbf{K}^{c})}$${\operatorname{Thom}(\operatorname{Spc}\mathbf{K}^{c})}$${\Theta_{\operatorname{supp}^{\rm B}}(S):=\{A\in\mathbf{K}^{c}:\operatorname{supp}^{\rm B}(A)\subseteq S\}}$${S}$
Proof.
See [Bal05] for the case that $\mathbf{K}^{c}$ is commutative, [MR23] for the case that $\mathbf{K}^{c}$ is completely prime, [Row24] for the case that $\mathbf{K}^{c}$ is $\operatorname{Spc}$-Noetherian, and [NVY23] for the case that $\mathbf{K}^{c}$ is finitely-generated (i.e., has a thick generator). ∎
In practice, it is difficult to determine the Balmer spectrum concretely. To do so, we turn to the theory abstract support varieties. A support datum $(X,\sigma)$ on $\mathbf{K}^{c}$ consists of a map $\sigma$ from objects of $\mathbf{K}^{c}$ to closed subsets of some $T_{0}$ topological space $X$, satisfying
  1. (1)
    $\sigma(\mathbf{1})=X,\sigma(0)=\varnothing$;
  2. (2)
    $\sigma(A\oplus B)=\sigma(A)\cup\sigma(B)$ for all $A,B\in\mathbf{K}^{c}$;
  3. (3)
    $\sigma(\Sigma A)=\sigma(A)$ for all $A\in\mathbf{K}^{c};$
  4. (4)
    if $A\to B\to C\to\Sigma A$ is a distinguished triangle in $\mathbf{K}^{c}$, then $\sigma(A)\subseteq\sigma(B)\cup\sigma(C);$
  5. (5)
    $\sigma(A\otimes B)\subseteq\sigma(A)\cap\sigma(B)$ for all $A,B\in\mathbf{K}^{c}$.
When $X$ is clear from context (or if the particular topological space $X$ is immaterial to the discussion at hand), we will refer to $\sigma$ as the support datum, or, alternatively, as the support map. We will call a support map $\sigma$ tensorial if it satisfies
  1. (6)
    $\bigcup_{B\in\mathbf{K}^{c}}\sigma(A\otimes B\otimes C)=\sigma(A)\cap\sigma(C)$, for all $A$ and $C$ in $\mathbf{K}^{c}$;
faithful if
  1. (7)
    $\sigma(A)=\varnothing\Rightarrow A\cong 0$, for all $A\in\mathbf{K}^{c}$;
and realizing if
  1. (8)
    for any closed subset $V$ of $X$, there exists an object $A$ with $\sigma(A)=V$.
Remark 2.3.
The assumption that the topological space $X$ be $T_{0}$ is not included in the axioms put forward in [Bal05], but we include it here since it will be necessary throughout the paper. Note that the Balmer spectrum $\operatorname{Spc}\mathbf{K}^{c}$ is automatically $T_{0}$.
Example 2.4.
The pair $(\operatorname{Spc}\mathbf{K}^{c},\operatorname{supp}^{\rm B})$ is a tensorial and faithful support datum. Recall that we are assuming throughout that $\mathbf{K}^{c}$ is rigid; this is necessary for the faithfulness of $\operatorname{supp}^{\rm B}$ ([NVY22b, Proposition 4.1.1]). If $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian, then $\operatorname{supp}^{\rm B}$ is also realizing, by Theorem 2.2.
Tensoriality is particularly important, due to the universal property of the Balmer spectrum:
Theorem 2.5 ([NVY22a, Theorem 1.3.1]).
If $(X,\sigma)$ is a tensorial support datum on $\mathbf{K}^{c}$, then there is a unique continuous map
[…]
such that $\sigma(A)=\eta^{-1}(\operatorname{supp}^{\rm B}(A))$ for all $A\in\mathbf{K}^{c}.$
The following result about the universal map $\eta$ is implied by the proof of [NVY22a, Theorem 6.2.1(b)(ii)].
Proposition 2.6.
Let $(X,\sigma)$ be a realizing support datum on a monoidal triangulated category $\mathbf{K}^{c}$. Then the universal map $\eta:X\to\operatorname{Spc}\mathbf{K}^{c}$ as in Theorem 2.5 is injective.
Note this proposition uses the fact that $X$ is assumed to be $T_{0}$, which is an assumption we make about all support data in this paper.
Notation 2.7.
Given a support datum $\sigma$, we have analogues of the maps from Theorem 2.2: given a thick ideal $\mathbf{I}$ of $\mathbf{K}^{c}$, and a subset $S$ of $X$, we will denote
$\Phi_{\sigma}(\mathbf{I})$ $:=\bigcup_{A\in\mathbf{I}}\sigma(A),\text{ and}$
$\Theta_{\sigma}(S)$ :=$:=\{A\in\mathbf{K}^{c}:\sigma(A)\subseteq S\}.$
When there is no risk of confusion, we will drop the subscript, and write $\Phi$ and $\Theta$ for $\Phi_{\sigma}$ and $\Theta_{\sigma}$, respectively.
This paper will be primarily concerned with extensions of a support $(X,\sigma)$ to the big category $\mathbf{K}$. By an extension, we mean the following.
Definition 2.8.
Let $(X,\sigma)$ be a support datum on $\mathbf{K}^{c}$. We say that $(X,\widetilde{\sigma})$ is an extension of $(X,\sigma)$ if $\widetilde{\sigma}$ is a map from $\mathbf{K}$ to (not necessarily closed) subsets of $X$, and
  1. (1)
    $\widetilde{\sigma}(A)=\sigma(A)$ for all $A$ compact;
  2. (2)
    $\widetilde{\sigma}(\coprod_{i\in I}A_{i})=\bigcup_{i\in I}\widetilde{\sigma}(A_{i})$ for any set-indexed collection $\{A_{i}|i\in I\}$ of objects of $\mathbf{K}$;
  3. (3)
    $\widetilde{\sigma}(\Sigma A)=\widetilde{\sigma}(A)$ for all $A\in\mathbf{K};$
  4. (4)
    if $A\to B\to C\to\Sigma A$ is a distinguished triangle in $\mathbf{K}$, then $\widetilde{\sigma}(A)\subseteq\widetilde{\sigma}(B)\cup\widetilde{\sigma}(C);$
  5. (5)
    $\widetilde{\sigma}(A\otimes B)\subseteq\widetilde{\sigma}(A)\cap\widetilde{\sigma}(B)$ for all $A,B\in\mathbf{K}$.
We call $\widetilde{\sigma}$ tensorial if
  1. (6)
    $\bigcup_{B\in\mathbf{K}^{c}}\widetilde{\sigma}(A\otimes B\otimes C)=\widetilde{\sigma}(A)\cap\widetilde{\sigma}(C)$, for all $A$ in $\mathbf{K}$ and $C\in\mathbf{K}^{c}$,
and faithful if
  1. (7)
    $\widetilde{\sigma}(A)=\varnothing\Rightarrow A\cong 0$, for all $A\in\mathbf{K}$.
The following theorem is our primary motivation in providing an extension for a support: it is a necessary condition, in many cases, for providing concrete realizations of the Balmer spectrum and Balmer support.
Theorem 2.9 ([NVY22a, Theorem 1.4.1]).
Let $(X,\sigma)$ be a tensorial, faithful, realizing support with value in a Zariski space $X$ on a small monoidal triangulated category $\mathbf{K}^{c}$. If $(X,\sigma)$ admits a tensorial, faithful extension $(X,\widetilde{\sigma})$, then there is a homeomorphism
$\eta:X\xrightarrow{\sim}\operatorname{Spc}\mathbf{K}^{c}$
such that $\eta^{-1}(\operatorname{supp}^{\rm B}(A))=\sigma(A)$ for all $A\in\mathbf{K}^{c}$.
For reference later, we note the elementary fact that extended supports are compatible with taking localizing subcategories in the following way:
Lemma 2.10.
Let $\sigma$ be a support datum on $\mathbf{K}^{c}$ with extension $\widetilde{\sigma}$. Let $\mathbf{I}$ be a thick ideal of $\mathbf{K}^{c}$, and let $\operatorname{Loc}(\mathbf{I})$ denote the smallest localizing subcategory of $\mathbf{K}$ containing $\mathbf{I}$. If $A$ is in $\operatorname{Loc}(\mathbf{I})$, then $\widetilde{\sigma}(A)\subseteq\Phi_{\sigma}(\mathbf{I}).$
Proof.
Using the axioms of extension, it is clear that the collection of objects
$\mathbf{L}:=\{B\in\mathbf{K}:\widetilde{\sigma}(B)\subseteq\Phi_{\sigma}(\mathbf{I})\}$
is a localizing subcategory. Since $\mathbf{I}\subseteq\mathbf{L}$ by the fact that $\sigma=\widetilde{\sigma}$ on compact objects, it follows that $\operatorname{Loc}(\mathbf{I})\subseteq\mathbf{L}$. ∎
3. Rickard idempotents and formal extensions of support
In this section we review background on Rickard idempotent functors, and subsequently construct formal extensions of support theories.
Let $\mathbf{K}$ be a big monoidal triangulated category, and let $\mathbf{I}$ be a thick ideal of $\mathbf{K}^{c}$. Then the Verdier quotient functor
$\mathbf{K}\to\mathbf{K}/\operatorname{Loc}(\mathbf{I})$
admits a right adjoint, by a general argument of Neeman (see [Nee92b, Lemma 1.7]) which was inspired by older arguments of Adams and Bousfield. By composing the quotient functor with its adjoint, one obtains a functor $L_{\mathbf{I}}:\mathbf{K}\to\mathbf{K}$ which takes value in $\operatorname{Loc}(\mathbf{I})^{\perp}$, the subcategory of $\mathbf{K}$ of objects $B$ such that $\operatorname{Hom}_{\mathbf{K}}(A,B)=0$ for all $A\in\operatorname{Loc}(\mathbf{I})$. By analogous argument, there exists a right adjoint to the inclusion functor
$\operatorname{Loc}(\mathbf{I})\to\mathbf{K},$
and again by composition we get a functor $\it{\Gamma}_{\mathbf{I}}:\mathbf{K}\to\mathbf{K}$ which takes value in $\operatorname{Loc}(\mathbf{I})$; see also [Kra10, §7.2]. A similar construction was given by Miller in [Mil92] for the category of spectra. Given an object $A$, the object $L_{\mathbf{I}}A$ can be described directly by a process of repeated homotopy colimits given by ranging over all possible maps from objects of $\mathbf{I}$. This construction was also used directly by Rickard in [Ric97], who subsequently used the functors $\it{\Gamma}_{\mathbf{I}}$ and $L_{\mathbf{I}}$ in joint work with Benson and Carlson to define the first extended supports in modular representation theory [BCR95, BCR96]. The functors $\it{\Gamma}_{\mathbf{I}}$ and $L_{\mathbf{I}}$, which are now commonly referred to as Rickard idempotent functors, satisfy the following universal property:
Theorem 3.1 ([Nee92b, Ric97]).
For any thick subcategory $\mathbf{I}$ of $\mathbf{K}^{c}$, there exist unique functors $L_{\mathbf{I}},{\it{\Gamma}}_{\mathbf{I}}:\mathbf{K}\to\mathbf{K}$ such that for each $A\in\mathbf{K}$, there is a distinguished triangle
${\it{\Gamma}}_{\mathbf{I}}A\to A\to L_{\mathbf{I}}A\to\Sigma{\it{\Gamma}}_{\mathbf{I}}A$
with ${\it{\Gamma}}_{\mathbf{I}}A\in\operatorname{Loc}(\mathbf{I})$ and $L_{\mathbf{I}}A\in\operatorname{Loc}(\mathbf{I})^{\perp}.$
We summarize some basic properties satisfied by the Rickard idempotent functors which we will use throughout.
Lemma 3.2.
Let $\mathbf{I}$ and $\mathbf{J}$ be thick ideals of $\mathbf{K}^{c}$ such that $\mathbf{J}\subseteq\mathbf{I}$, and let $A\in\mathbf{K}$. Then:
  1. (1)
    ${\it{\Gamma}}_{\mathbf{I}}A\cong A$ if and only if $L_{\mathbf{I}}A\cong 0$ if and only if $A\in\operatorname{Loc}(\mathbf{I})$;
  2. (2)
    $A\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\cong{\it{\Gamma}}_{\mathbf{I}}A\cong{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\otimes A;$
  3. (3)
    $A\otimes L_{\mathbf{I}}\mathbf{1}\cong L_{\mathbf{I}}A\cong L_{\mathbf{I}}\mathbf{1}\otimes A;$
  4. (4)
    $L_{\mathbf{I}}\mathbf{1}\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\cong 0;$
  5. (5)
    ${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\otimes{\it{\Gamma}}_{\mathbf{J}}\mathbf{1}\cong{\it{\Gamma}}_{\mathbf{J}}\mathbf{1}$ and $L_{\mathbf{I}}\mathbf{1}\otimes L_{\mathbf{J}}\mathbf{1}\cong L_{\mathbf{I}}\mathbf{1};$
  6. (6)
    there is a unique morphism of triangles
    ${{\it{\Gamma}}_{\mathbf{J}}\mathbf{1}}$${\mathbf{1}}$${L_{\mathbf{J}}\mathbf{1}}$${{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}}$${\mathbf{1}}$${L_{\mathbf{I}}\mathbf{1}}$
Proof.
(1) is clear. (2) and (3) follow from [NVY23, Lemma 3.2.1]. (4) is a direct consequence of (1), (2), and the fact that ${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\in\operatorname{Loc}(\mathbf{I})$. (5) is straightforward. For (6), see [BF11, Remark 2.9].∎
Note in particular that by Lemma 3.2(2) and (3), both $L_{\mathbf{I}}\mathbf{1}$ and ${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}$ tensor-commute with all objects of $\mathbf{K}$. This fact will be used frequently throughout the paper.
Notation 3.3.
Recall that given a support datum $(X,\sigma)$ and a specialization-closed subset $S$ of $X$, we have the thick ideal $\Theta(S)$ defined in Notation 2.7. To avoid burdening the notation, we set
${\it{\Gamma}}_{S}:={\it{\Gamma}}_{\Theta(S)},\;\;L_{S}:=L_{\Theta(S)}.$
We now define extended support for an arbitrary support datum $(\sigma,X)$; this follows the definitions put forward by Benson–Iyengar–Krause [BIK08] and Balmer–Favi [BF11] from the commutative case.
Let $x$ be a point of $X$. We set
$\mathcal{Z}(x)$ :=$:=\{y\in X:x\not\in\overline{\{y\}}\},$
$\mathcal{V}(x)$ :=$:=\overline{\{x\}}.$
Note that $\mathcal{Z}(x)$ and $\mathcal{V}(x)$ are both specialization-closed, and, since we assume that $X$ is $T_{0}$, we have $\{x\}=\mathcal{V}(x)\backslash(\mathcal{V}(x)\cap\mathcal{Z}(x))$. We now set
${\it{\Gamma}}_{x}:={\it{\Gamma}}_{\mathcal{V}(x)}\otimes L_{\mathcal{Z}(x)}.$
Remark 3.4.
The results in this paper will focus on support data based on Noetherian spaces $X$. For non-Noetherian spaces, one can define $\it{\Gamma}_{x}$ in a more general way using Thomason subsets, as discussed in [Ste14a, §2.1], to get around the issue that $\mathcal{V}(x)$ might not be Thomason. Recall that in a Noetherian space, any specialization closed subset is automatically a Thomason subset.
Remark 3.5.
We caution the reader that if $x$ is a point, then ${\it{\Gamma}}_{x}$ is not the same as ${\it{\Gamma}}_{\overline{\{x\}}}$, which is defined according to Notation 3.3.
Notation 3.6.
Unfortunately, the notation ${\it{\Gamma}}_{\mathbf{P}}$, for $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$, is now overloaded: (1) if we consider $\mathbf{P}$ as a thick ideal, it has a meaning as in Theorem 3.1, and, on the other hand, (2) if we consider $\mathbf{P}$ as a point of $\operatorname{Spc}\mathbf{K}^{c}$ equipped with the Balmer support $\operatorname{supp}^{\rm B}$, then we have
${\it{\Gamma}}_{\mathbf{P}}:={\it{\Gamma}}_{\mathcal{V}(\mathbf{P})}\otimes L_{\mathcal{Z}(\mathbf{P})}$
as defined above. For the remainder of the paper, if we write ${\it{\Gamma}}_{\mathbf{P}}$ for $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$, we will always mean (2), that is, we consider $\mathbf{P}$ as a point of $\operatorname{Spc}\mathbf{K}^{c}$ rather than as a thick ideal.
Definition 3.7.
Let $(X,\sigma)$ be a support datum on $\mathbf{K}^{c}$. Define the extension of $\sigma$, denoted $(X,\widetilde{\sigma})$, by
$\widetilde{\sigma}(A):=\{x\in X:{\it{\Gamma}}_{x}A\not=0\}$
for all $A\in\mathbf{K}.$
It is straightforward to check the following:
Proposition 3.8.
Let $(X,\sigma)$ be a tensorial realizing support on a small monoidal triangulated category $\mathbf{K}^{c}$. Then $(X,\widetilde{\sigma})$ is an extension of $(X,\sigma)$ (recall Definition 2.8).
Proof.
Since both functors ${\it{\Gamma}}_{\mathbf{I}}$ and $L_{\mathbf{I}}$ are monoidal triangulated functors commuting with coproducts, conditions (2)-(5) of Definition 2.8 are automatic. It remains to show that $\sigma(A)=\widetilde{\sigma}(A)$ for all compact objects $A$; the proof of this fact will proceed similarly to [BF11, Proposition 7.17], which proves the analogous result for $\sigma=\operatorname{supp}^{\rm B}$ in the commutative case. We need to show that for any compact $A$, we have $x\not\in\sigma(A)$ if and only if ${\it{\Gamma}}_{x}\mathbf{1}\otimes A\cong 0$.
Suppose $x\not\in\sigma(A)$; this is equivalent to $\sigma(A)\subseteq\mathcal{Z}(x)$. That in turn is equivalent to $L_{\mathcal{Z}(x)}\mathbf{1}\otimes A\cong 0$ by Lemma 3.2(1), which then immediately implies ${\it{\Gamma}}_{x}\mathbf{1}\otimes A\cong 0$.
On the other hand, assume ${\it{\Gamma}}_{x}\mathbf{1}\otimes A\cong 0$. By the assumption that $\sigma$ is realizing, we can find an object $C\in\mathbf{K}^{c}$ with $\sigma(C)=\mathcal{V}(x)$. Note that $C\otimes{\it{\Gamma}}_{\mathcal{V}(x)}\cong C$. For any $B\in\mathbf{K}^{c}$, we then find that
00 0 $\cong A\otimes{\it{\Gamma}}_{x}\mathbf{1}\otimes B\otimes C$
$\cong A\otimes L_{\mathcal{Z}(x)}\otimes B\otimes C$
[…]
$A\otimes B\otimes C$ $\in\operatorname{Loc}(\Theta(\mathcal{Z}(x)))$
[…]
$A\otimes B\otimes C$ $\in\Theta(\mathcal{Z}(x))$
[…]
$\sigma(A\otimes B\otimes C)$ $\subseteq\mathcal{Z}(x).$
Since we assume that $\sigma$ is tensorial, this implies that
$\bigcup_{B\in\mathbf{K}^{c}}\sigma(A\otimes B\otimes C)=\sigma(A)\cap\sigma(C)\subseteq\mathcal{Z}(x).$
Now note that $x\not\in\mathcal{Z}(x)$, but $x\in\mathcal{V}(x)$, and recall that we have assumed $\sigma(C)=\mathcal{V}(x)$. It follows that $x\not\in\sigma(A)$. ∎
4. Tensoriality of extended Balmer support
In this section, we prove that when $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian, the extension of the Balmer support is tensorial. This will follow closely the proof given in the commutative case by Balmer and Favi [BF11, Section 7].
We first set some notation.
Notation 4.1.
The extension of the Balmer support $\operatorname{supp}^{\rm B}$ as in Definition 3.7 will be denoted $\operatorname{Supp}^{\rm B}$.
To prove tensoriality, we first describe the extended Balmer supports of the Rickard idempotents.
Lemma 4.2.
Let $\mathbf{K}^{c}$ be a small $\operatorname{Spc}$-Noetherian monoidal triangulated category. Let $\mathbf{I}$ be any thick ideal of $\mathbf{K}^{c}$. Then $\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}\mathbf{1})=\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})$ and $\operatorname{Supp}^{\rm B}(L_{\mathbf{I}}\mathbf{1})=\operatorname{Spc}\mathbf{K}^{c}\backslash\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I}).$
Proof.
Let $A\in\mathbf{I}$, and let $\mathbf{P}\in\operatorname{supp}^{\rm B}(A).$ Then $A\otimes{\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\not=0.$ Since ${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\otimes A\cong A\cong A\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}$ by Lemma 3.2(1) and (2), it follows that ${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\otimes{\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\not=0,$ that is, $\mathbf{P}\in\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}\mathbf{1}).$ Hence
$\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})\subseteq\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}(A)).$
On the other hand, since ${\it{\Gamma}}_{\mathbf{I}}(A)$ is in $\operatorname{Loc}(\mathbf{I})$, we have $\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}(A))\subseteq\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})$ by Lemma 2.10.
Suppose $\mathbf{P}\in\operatorname{Supp}^{\rm B}(L_{\mathbf{I}}\mathbf{1})$, i.e. ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes L_{\mathbf{I}}\mathbf{1}\not=0.$ Recall that by 3.2, if $\Theta(\mathcal{V}(\mathbf{P}))\subseteq\mathbf{I}$, this would imply that ${\it{\Gamma}}_{\mathcal{V}(\mathbf{P})}\mathbf{1}\otimes L_{\mathbf{I}}\mathbf{1}\cong 0$, which would immediately imply that ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes L_{\mathbf{I}}\mathbf{1}\cong 0$, a contradiction, so we have $\Theta(\mathcal{V}(\mathbf{P}))\not\subseteq\mathbf{I}$. Using the Noetherianity of $\operatorname{Spc}\mathbf{K}^{c}$, by Theorem 2.2 we have
$\mathcal{V}(\mathbf{P})\not\subseteq\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I}),$
i.e. $\mathbf{P}\not\in\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})$. It follows that
$\operatorname{Supp}^{\rm B}(L_{\mathbf{I}}\mathbf{1})\subseteq\operatorname{Spc}\mathbf{K}^{c}\backslash\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I}).$
Since we have a distinguished triangle
${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\to\mathbf{1}\to L_{\mathbf{I}}\mathbf{1},$
we have that $\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}\mathbf{1})\cup\operatorname{Supp}^{\rm B}(L_{\mathbf{I}}\mathbf{1})=\operatorname{Spc}\mathbf{K}^{c},$ and so the opposite containment
$\operatorname{Supp}^{\rm B}(L_{\mathbf{I}}\mathbf{1})\supseteq\operatorname{Spc}\mathbf{K}^{c}\backslash\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})$
holds as well. ∎
We now prove that $\operatorname{Supp}^{\rm B}$ satisfies a tensor product property when tensoring by Rickard idempotents.
Lemma 4.3.
Let $\mathbf{K}^{c}$ be a small $\operatorname{Spc}$-Noetherian monoidal triangulated category. Let $A$ be an object of $\mathbf{K}$, and let $\mathbf{I}$ be a thick ideal of $\mathbf{K}^{c}$. Then
$\operatorname{Supp}^{\rm B}(A\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1})$ $=\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}\mathbf{1})=\operatorname{Supp}^{\rm B}(A)\cap\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I}),$
$\operatorname{Supp}^{\rm B}(A\otimes L_{\mathbf{I}}\mathbf{1})$ $=\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}(L_{\mathbf{I}}\mathbf{1})=\operatorname{Supp}^{\rm B}(A)\cap(\operatorname{Spc}\mathbf{K}^{c}\backslash\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})).$
Proof.
Since $\operatorname{Supp}^{\rm B}$ is an extended support, the containments
$\operatorname{Supp}^{\rm B}(A\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1})$ $\subseteq\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}\mathbf{1}),$
$\operatorname{Supp}^{\rm B}(A\otimes L_{\mathbf{I}}\mathbf{1})$ $\subseteq\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}(L_{\mathbf{I}}\mathbf{1})$
are automatic. For the opposite containment, suppose $\mathbf{P}\in\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathbf{I}}\mathbf{1})$. In particular, by Lemma 4.2, $\mathbf{P}\in\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})$, in which case $\mathcal{V}(\mathbf{P})\subseteq\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})$, and hence $\Theta(\mathcal{V}(\mathbf{P}))\subseteq\mathbf{I}$. It follows from Lemma 3.2(5) that ${\it{\Gamma}}_{\mathcal{V}(\mathbf{P})}\mathbf{1}\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\cong{\it{\Gamma}}_{\mathcal{V}(\mathbf{P})}\mathbf{1}$, and so
$A\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\otimes{\it{\Gamma}}_{\mathbf{P}}\mathbf{1}$ $\cong A\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\otimes{\it{\Gamma}}_{\mathcal{V}(\mathbf{P})}\mathbf{1}\otimes L_{\mathcal{Z}(\mathbf{P})}\mathbf{1}$
$\cong A\otimes{\it{\Gamma}}_{\mathcal{V}(\mathbf{P})}\mathbf{1}\otimes L_{\mathcal{Z}(\mathbf{P})}\mathbf{1}$
$\cong A\otimes{\it{\Gamma}}_{\mathbf{P}}\mathbf{1}$
≇$\not\cong 0,$
so $\mathbf{P}\in\operatorname{Supp}^{\rm B}(A\otimes{\it{\Gamma}}_{\mathbf{I}}\mathbf{1})$. The second claim is similar. ∎
We are now able to prove tensoriality of the Balmer spectrum by adapting the proof from the commutative case, see [BF11, Theorem 7.22].
Proposition 4.4.
Let $\mathbf{K}^{c}$ be a small $\operatorname{Spc}$-Noetherian monoidal triangulated category. Then the extended Balmer support $\operatorname{Supp}^{\rm B}$ is tensorial.
Proof.
Let $A\in\mathbf{K}$ and $C\in\mathbf{K}^{c}$. Since $\operatorname{Supp}^{\rm B}$ is an extended support, it is automatic that
$\bigcup_{B\in\mathbf{K}^{c}}\operatorname{Supp}^{\rm B}(A\otimes B\otimes C)\subseteq\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}(C).$
We must show the opposite containment. Set $V:=\operatorname{supp}^{\rm B}(C)$. We claim that $C$ and ${\it{\Gamma}}_{V}\mathbf{1}$ generate the same localizing $\otimes$-ideal. Indeed, ${\it{\Gamma}}_{V}\mathbf{1}\otimes C\cong C$ by Lemma 3.2 since $C\in\Theta(V)$, so $C$ is in the localizing $\otimes$-ideal generated by ${\it{\Gamma}}_{V}\mathbf{1}$; and on the other hand, ${\it{\Gamma}}_{V}\mathbf{1}$ is by definition in $\operatorname{Loc}(\Theta(V))$, which is equal to $\operatorname{Loc}(\langle C\rangle)$ by Theorem 2.2, so ${\it{\Gamma}}_{V}\mathbf{1}$ is in the localizing $\otimes$-ideal generated by $C$. Since the kernel of ${\it{\Gamma}}_{\mathbf{P}}\otimes-$ is a localizing tensor-ideal, it follows that $\operatorname{Supp}^{\rm B}(C)=\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{V}\mathbf{1})$. Hence
$\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}(C)=\operatorname{Supp}^{\rm B}(A)\cap\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{V}\mathbf{1})=\operatorname{Supp}^{\rm B}(A\otimes{\it{\Gamma}}_{V}\mathbf{1})$
by Lemma 4.3. Suppose $\mathbf{P}\in\operatorname{Supp}^{\rm B}(A\otimes{\it{\Gamma}}_{V}\mathbf{1})$, so that ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes A\otimes{\it{\Gamma}}_{V}\mathbf{1}\not=0$. Since ${\it{\Gamma}}_{V}\mathbf{1}\in\operatorname{Loc}(\langle C\rangle)$, it is straightforward to check that there must be some $D\in\langle C\rangle$ with ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes A\otimes D\not=0$. This implies there is some $B\in\mathbf{K}^{c}$ with ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes A\otimes B\otimes C\not=0$ by the following argument. Suppose ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes A\otimes B\otimes C=0$ for all $B\in\mathbf{K}^{c}$. Consider
$\mathbf{J}:=\{E\in\mathbf{K}^{c}:\;\;{\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes A\otimes B\otimes E=0\;\;\forall B\in\mathbf{K}^{c}\}.$
It can be readily checked that $\mathbf{J}$ is a thick ideal, and since it contains $C$, it must contain all of $\langle C\rangle$, a contradiction. Hence there exists $B$ with ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\otimes A\otimes B\otimes C\not=0,$ and we have proven the reverse containment.
5. Faithfulness of extended comparative support
Throughout this section, we will assume that
[…]
the reason being that either allows us to apply Theorem 2.2. We will assume that $(X,\sigma)$ is a tensorial support such that there exists a map $\rho:\operatorname{Spc}\mathbf{K}^{c}\to X$ satisfying
  1. (1)
    $\rho\circ\eta=\operatorname{id}_{X}$, and
  2. (2)
    $(\eta\circ\rho)(\mathbf{P})\subseteq\mathbf{P}$ for all $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$,
where $\eta:X\to\operatorname{Spc}\mathbf{K}^{c}$ is the universal map that exists by Theorem 2.5. Recall that by the definition of the topology on $\operatorname{Spc}\mathbf{K}^{c}$, (2) is equivalent to the condition that $\eta\rho(\mathbf{P})\in\overline{\{\mathbf{P}\}}.$
Definition 5.1.
A support datum $(X,\sigma)$ is called comparative if it satisfies (1) and (2) above for some map $\rho$. The map $\rho$ will be called a comparison map.
The terminology is intended to reflect that $\rho$ is an analogue of the comparison map (also denoted $\rho$) produced by Balmer in [Bal10, Theorem 5.3]. In this section, we will show that the extension $\widetilde{\sigma}$ of a comparative support $\sigma$ to $\mathbf{K}^{c}$ is faithful, as long as $X$ is a Zariski space and $\mathbf{K}^{c}$ is finitely-generated (recall, this means that there exists an object $G$ in $\mathbf{K}^{c}$ such that $\mathbf{K}^{c}$ is the smallest thick subcategory of $\mathbf{K}^{c}$ containing $G$) or $\operatorname{Spc}$-Noetherian. Recall here that a sober topological space is one in which every irreducible closed set has a unique generic point, and a Zariski space is one which is Noetherian and sober. The main applications of this theorem are twofold: (1) for $\sigma=\operatorname{supp}^{\rm B}$, which is manifestly comparative (both $\rho$ and $\eta$ are the identity maps), and (2) the central cohomological support for the stable category of a finite tensor category, a version of support introduced in [NVY24], which we explore in Section 8.
We begin by noting a few elementary properties of comparative support that will be useful.
Lemma 5.2.
Let $(X,\sigma)$ be a comparative support on $\mathbf{K}^{c}$, and let $X$ be sober. Then
  1. (1)
    the comparison map $\rho$ is a closed map;
  2. (2)
    $\sigma(A)=\rho(\operatorname{supp}^{\rm B}(A));$
  3. (3)
    $\sigma$ is faithful;
  4. (4)
    if $V\subseteq X$ is a Thomason-closed subset, then $\rho^{-1}(V)$ is Thomason-closed in $\operatorname{Spc}\mathbf{K}^{c}$;
  5. (5)
    if $X$ is Noetherian and $\mathbf{K}^{c}$ is finitely-generated, then $\sigma$ is realizing.
Proof.
(1): let $V$ be a closed subset of $\operatorname{Spc}\mathbf{K}^{c}$, and let $x$ be in the closure of $\rho(V)$. Then there exists some $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$ such that $x$ is in the closure of $\rho(\mathbf{P})$, so $\eta(x)$ is in the closure of $\{\eta\rho(\mathbf{P})\}$. Since $\eta\rho(\mathbf{P})\subseteq\mathbf{P}$, we also have $\eta\rho(\mathbf{P})$ is in the closure of $\{\mathbf{P}\}$, so $\eta\rho(\mathbf{P})$ is in $V$. It follows that $\eta(x)\in V$, so $x=\rho\eta(x)$ is in $\rho(V)$.
(2): recall that by the universal property of $\operatorname{Spc}\mathbf{K}^{c}$, we have $\sigma(A)=\eta^{-1}(\operatorname{supp}^{\rm B}(A))$. Suppose $x\in\rho(\operatorname{supp}^{\rm B}(A))$, that is, there exists $\mathbf{P}\in\operatorname{supp}^{\rm B}(A)$ such that $\rho(\mathbf{P})=x.$ Then since $\eta\rho(\mathbf{P})=\eta(x)\subseteq\mathbf{P}$, and since $\mathbf{P}\in\operatorname{supp}^{\rm B}(A)$, it follows that $\eta(x)\in\operatorname{supp}^{\rm B}(A)$, and so $x\in\eta^{-1}(\operatorname{supp}^{\rm B}(A))$. For the other direction, if $\eta(x)\in\operatorname{supp}^{\rm B}(A)$ for some point $x$, then since $x=\rho\eta(x),$ we have $x\in\rho(\operatorname{supp}^{\rm B}(A))$.
(3): if $A\in\mathbf{K}^{c}$ with $\sigma(A)=\varnothing,$ then (2) implies $\operatorname{supp}^{\rm B}(A)=\varnothing$, which by faithfulness of $\operatorname{supp}^{\rm B}$ implies that $A\cong 0$.
(4): suppose $V$ is Thomason-closed, that is, $V=\bigcup_{k\in K}V_{k}$ where $V_{k}$ is closed and $X\backslash V_{k}$ is quasicompact, for each $k$. Then $\rho^{-1}(V)=\bigcup_{k\in K}\rho^{-1}(V_{k}),$ so it suffices to show that each $\rho^{-1}(V_{k})$ has quasicompact complement. Since $\operatorname{Spc}\mathbf{K}^{c}\backslash\rho^{-1}(V_{k})=\rho^{-1}(X\backslash V_{k})$, it suffices to prove that the inverse image of a quasicompact open set is quasicompact.
Suppose $U$ is a quasicompact open subset of $X$, and suppose $\rho^{-1}(U)$ has an open cover $\{U_{i}\}_{i\in I}$. We must show that it admits some finite subcover. We claim that the collection of open sets
$\{X\backslash(\rho(\operatorname{Spc}\mathbf{K}^{c}\backslash U_{i}))\}_{i\in I}$
forms an open cover of $U$ (each of these sets is open since $\rho$ is a closed map by (1)). Indeed, if $x\in U$, then $\eta(x)\in\rho^{-1}(U)$, so $\eta(x)\in U_{i}$ for some $i$; we claim that $x\in X\backslash(\rho(\operatorname{Spc}\mathbf{K}^{c}\backslash U_{i}))$ for this particular $i$. Suppose to the contrary. Then $x\in\rho(\operatorname{Spc}\mathbf{K}^{c}\backslash U_{i})$, so there exists $\mathbf{P}\not\in U_{i}$ with $x=\rho(\mathbf{P})$. But $\eta(x)=\eta\rho(\mathbf{P})\subseteq\mathbf{P}$, so any open set containing $\eta(x)$ must contain $\mathbf{P}$, and $\eta(x)\in U_{i}$, so $\mathbf{P}\in U_{i}$, a contradiction. Our claim that $x\in X\backslash(\rho(\operatorname{Spc}\mathbf{K}^{c}\backslash U_{i}))$ follows.
Now, since $U$ is finite, there exists a finite subset $J\subseteq I$ such that
$\{X\backslash(\rho(\operatorname{Spc}\mathbf{K}^{c}\backslash U_{j}))\}_{j\in J}$
covers $U$. We now claim that $\{U_{j}\}_{j\in J}$ covers $\rho^{-1}(U)$, at which point our proof will be complete. It is enough to check that each $\eta(x),$ for $x\in U$, is in some $U_{j}$ for $j\in J$, since again if $\mathbf{P}\in\rho^{-1}(U)$ then $\mathbf{P}$ is in any open subset containing $\eta\rho(\mathbf{P})$. Since $x\in U$, it is in some $X\backslash(\rho(\operatorname{Spc}\mathbf{K}^{c}\backslash U_{j}))$ for some $j\in J$; we claim that $\eta(x)\in U_{j}$. Suppose to the contrary. Then $\eta(x)\in\operatorname{Spc}\mathbf{K}^{c}\backslash U_{j}$, and so $x=\rho\eta(x)\in\rho(\operatorname{Spc}\mathbf{K}^{c}\backslash U_{j})$, a contradiction. Hence $\eta(x)\in U_{j}$, and the collection
$\{U_{j}\}_{j\in J}$
is indeed a finite cover of $\rho^{-1}(U)$, completing the proof.
(5) If $V\subseteq X$ is a closed subset, then it is Thomason-closed, since $X$ is Noetherian. By (4), $\rho^{-1}(V)$ is Thomason-closed in $\operatorname{Spc}\mathbf{K}^{c}$, and so by Theorem 2.2 we have $\rho^{-1}(V)=\operatorname{supp}^{\rm B}(A)$ for some $A\in\mathbf{K}^{c}$, using the assumption that $\mathbf{K}^{c}$ is finitely-generated. But then $\sigma(A)=\rho(\operatorname{supp}^{B}(A))=\rho(\rho^{-1}(V))=V$, since $\rho$ is surjective. ∎
Using this, we have an alternative characterization of $\sigma$ being comparative.
Proposition 5.3.
A realizing tensorial support $\sigma$ is comparative if and only if for all $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$, there exists a unique generic point of $\eta^{-1}(\overline{\{\mathbf{P}\}})$.
Proof.
Note that the condition is clear if $\mathbf{P}$ is in the image of $\eta$. Let us prove the $\Leftarrow$ implication first. Recall that by Proposition 2.6, since $\sigma$ is realizing, we have $\eta$ is injective. We construct $\rho$ by setting $\rho(\mathbf{P}):=x$, where $x$ is the unique generic point of $\rho^{-1}(\overline{\{\mathbf{P}\}})$. The map $\rho$ is continuous: any closed subset of $X$ may be written as $\sigma(A)$ for some $A\in\mathbf{K}^{c}$ by the realization assumption, and then we may verify that
$\rho^{-1}(\sigma(A))$ $=\rho^{-1}(\eta^{-1}(\operatorname{supp}^{\rm B}(A)))$
=$=\{\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}:\rho(\mathbf{P})\in\eta^{-1}(\operatorname{supp}^{\rm B}(A))\}$
=$=\{\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}:\eta(\rho(\mathbf{P}))\in\operatorname{supp}^{\rm B}(A)\}$
$=\operatorname{supp}^{\rm B}(A).$
It is clear that $\rho\circ\eta=\operatorname{id}_{X}$, and so if $\mathbf{P}$ is in the image of $\eta$ then $\eta\rho(\mathbf{P})=\eta\rho\eta(x)=\eta(x)=\mathbf{P}$ for some $x$. On the other hand, if $\mathbf{P}$ is not in the image of $\eta$, then since $\rho(\mathbf{P})\in\eta^{-1}(\overline{\{\mathbf{P}\}})$ by the definition of $\rho$, we have $\eta\rho(\mathbf{P})\subseteq\mathbf{P}$, and so $\sigma$ is comparative.
Let us now prove the $\Rightarrow$ direction. We claim that $\rho(\mathbf{P})$ is the unique generic point of $\eta^{-1}(\overline{\{\mathbf{P}\}})$, i.e. that $\overline{\rho(\mathbf{P})}=\eta^{-1}(\overline{\{\mathbf{P}\}})$. First, by the comparative hypothesis, $\eta(\rho(\mathbf{P}))\in\overline{\{\mathbf{P}\}}$, so $\rho(\mathbf{P})\in\eta^{-1}(\overline{\{\mathbf{P}\}})$. But this is a closed set, so $\overline{\rho(\mathbf{P})}\subseteq\eta^{-1}(\overline{\{\mathbf{P}\}})$. It remains only to prove the reverse inclusion. Take any $x\in\eta^{-1}(\overline{\{\mathbf{P}\}})$, hence $\eta(x)\in\overline{\{\mathbf{P}\}}$. Since $\rho$ is closed by Lemma 5.2(1),
$x=\rho\circ\eta(x)\in\rho(\overline{\{\mathbf{P}\}})=\overline{\rho(\mathbf{P})}\implies\eta^{-1}(\overline{\{\mathbf{P}\}})\subseteq\overline{\rho(\mathbf{P})}.$
Remark 5.4.
In fact, this proof implies that if $\sigma$ is realizing and comparative, then the comparison map $\rho$ is uniquely constructed from $\eta:X\hookrightarrow\operatorname{Spc}\mathbf{K}^{c}$, and $\rho(\mathbf{P})$ is the unique generic point of $\eta^{-1}(\overline{\{\mathbf{P}\}})$.
The next lemma is the primary tool which allows us to transport arguments using the bijection between Thomason-closed subsets of $\operatorname{Spc}\mathbf{K}^{c}$ and thick ideals of $\mathbf{K}^{c}$ to arguments based on the support $(X,\sigma).$
Lemma 5.5.
Let $\sigma$ be a comparative support with value in a Zariski space $X$ on a small monoidal triangulated category $\mathbf{K}^{c}$ satisfying (5.1). Let $V$ and $W$ be specialization-closed subsets of $X$. We have $\langle\Theta(V)\cup\Theta(W)\rangle=\Theta(V\cup W)$.
Recall that $\Theta(S)$ is the thick ideal $\{A\in\mathbf{K}^{c}:\sigma(A)\subseteq S\}$ for any specialization-closed subset $S$ of $X$.
Proof.
It is clear by definition that $\langle\Theta(V)\cup\Theta(W)\rangle\subseteq\Theta(V\cup W)$. To show the other direction, suppose $\sigma(A)\subseteq V\cup W$. Since $\sigma(A)=\rho(\operatorname{supp}^{\rm B}(A))$ by Lemma 5.2(2), $\operatorname{supp}^{\rm B}(A)\subseteq\rho^{-1}(V)\cup\rho^{-1}(W)$. Since $V$ and $W$ are Thomason-closed (indeed, every specialization-closed subset of $X$ is Thomason-closed, since $X$ is Noetherian), we know that $\rho^{-1}(V)$ and $\rho^{-1}(W)$ are Thomason-closed, by Lemma 5.2(4). We also know that $\operatorname{supp}^{\rm B}(A)$ is Thomason-closed, by Theorem 2.2. It follows that $V^{\prime}:=\operatorname{supp}^{\rm B}(A)\cap\rho^{-1}(V)$ and $W^{\prime}:=\operatorname{supp}^{\rm B}(A)\cap\rho^{-1}(W)$ are Thomason-closed, and by Theorem 2.2, there are thick ideals $\mathbf{I}$ and $\mathbf{J}$ with
$\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{I})=V^{\prime},\;\;\;\Phi_{\operatorname{supp}^{\rm B}}(\mathbf{J})=W^{\prime}.$
Again by Theorem 2.2, since $\operatorname{supp}^{\rm B}(A)\subseteq V^{\prime}\cup W^{\prime}$, we have $A\in\langle\mathbf{I},\mathbf{J}\rangle$. But now note that since $\operatorname{supp}^{\rm B}(B)\subseteq V^{\prime}\subseteq\rho^{-1}(V)$ for all $B\in\mathbf{I}$, and $\operatorname{supp}^{\rm B}(C)\subseteq W^{\prime}\subseteq\rho^{-1}(W)$ for all $C\in\mathbf{J}$, we have that $\mathbf{I}\subseteq\Theta(V)$ and $\mathbf{J}\subseteq\Theta(W)$. That is, we have
$A\in\langle\mathbf{I},\mathbf{J}\rangle\subseteq\langle\Theta(V),\Theta(W)\rangle,$
and the proof is complete.
We now leverage Lemma 5.5 to prove several useful identities for Rickard idempotents associated to specialization-closed subsets of $X$.
Corollary 5.6.
Let $\sigma$ be a comparative support with value in a Zariski space $X$ on a small monoidal triangulated category $\mathbf{K}^{c}$ satisfying (5.1). Let $V$ and $W$ be specialization-closed subsets of $X$. Then $L_{V\cup W}\mathbf{1}\cong L_{V}L_{W}\mathbf{1}.$
Proof.
Applying $L_{V}L_{W}$ to the distinguished triangle
${\it{\Gamma}}_{V\cup W}\mathbf{1}\to\mathbf{1}\to L_{V\cup W}\mathbf{1},$
we get
$L_{V}L_{W}{\it{\Gamma}}_{V\cup W}\mathbf{1}\to L_{V}L_{W}\mathbf{1}\to L_{V\cup W}\mathbf{1},$
by Lemma 3.2. But $L_{V}L_{W}A\in\langle\Theta(V)\cup\Theta(W)\rangle^{\perp}$ for any object $A$, by [NVY23, Lemma 3.2.2(a)]. Now note that $L_{V}L_{W}{\it{\Gamma}}_{V\cup W}\mathbf{1}\in\Theta(V\cup W)=\langle\Theta(V)\cup\Theta(W)\rangle$ by Lemma 5.5, and it follows that $L_{V}L_{W}{\it{\Gamma}}_{V\cup W}\mathbf{1}\cong 0$, hence $L_{V}L_{W}\mathbf{1}\cong L_{V\cup W}\mathbf{1}$. ∎
Lemma 5.7.
Let $\sigma$ be a comparative support with value in a Zariski space $X$ on a small monoidal triangulated category $\mathbf{K}^{c}$ satisfying (5.1). Suppose $Y,Z$ are specialization-closed subsets of $X$. Then $L_{Z}\mathbf{1}\otimes{\it{\Gamma}}_{Y}\mathbf{1}$ depends only on $Y\cap Z^{c}$.
Proof.
The proof is identical to [BF11, Lemma 7.4]; the main technical difficulty in adapting their proof, which is for the Balmer support, to an arbitrary comparative support is Lemma 5.5 above. We include the argument for the convenience of the reader. It suffices to show that $L_{Z}{\it{\Gamma}}_{Y}\mathbf{1}\cong L_{Z}{\it{\Gamma}}_{Y^{\prime}}\mathbf{1}$ for $Y\cap Z^{c}=Y^{\prime}\cap Z^{c}$ and $L_{Z}{\it{\Gamma}}_{Y}\mathbf{1}\cong L_{Z^{\prime}}{\it{\Gamma}}_{Y}\mathbf{1}$ for $Y\cap Z^{c}=Y\cap{Z^{\prime}}^{c}$, since in this case if $Y\cap Z^{c}=Y^{\prime}\cap Z^{\prime c}$ we can check
$L_{Z}{\it{\Gamma}}_{Y}\mathbf{1}\cong L_{Z}{\it{\Gamma}}_{Y\cap Y^{\prime}}\mathbf{1}\cong L_{Z\cup Z^{\prime}}{\it{\Gamma}}_{Y\cap Y^{\prime}}\mathbf{1}\cong L_{Z^{\prime}}{\it{\Gamma}}_{Y^{\prime}}\mathbf{1}.$
The two statements are proved the same way, so we will just show the first. In fact, it is enough to show that $L_{Z}{\it{\Gamma}}_{Y}\mathbf{1}\cong L_{Z}{\it{\Gamma}}_{Y^{\prime}}\mathbf{1}$ for $Y\subseteq Y^{\prime}$. In this case, we have a map of exact triangles
${{{\it{\Gamma}}_{Y}\mathbf{1}}}$${\mathbf{1}}$${{L_{Y}\mathbf{1}}}$${{{\it{\Gamma}}_{Y’}\mathbf{1}}}$${\mathbf{1}}$${{L_{Y’}\mathbf{1}}}$
by Lemma 3.2(6). Note that $Y\cup Z=(Y\cap Z^{c})\cup Z=(Y^{\prime}\cap{Z}^{c})\cup Z=Y^{\prime}\cup Z$, so Corollary 5.6 implies that we have an isomorphism
$L_{Y}\mathbf{1}\otimes L_{Z}\mathbf{1}\cong L_{Y\cup Z}\mathbf{1}\cong L_{Y^{\prime}\cup Z}\mathbf{1}\cong L_{Y^{\prime}}\mathbf{1}\otimes L_{Z}\mathbf{1}.$
This means that when we apply the exact functor $-\otimes L_{Z}\mathbf{1}$ to the triangles above, the second and third vertical maps become isomorphisms, hence ${\it{\Gamma}}_{Y}\mathbf{1}\otimes L_{Z}\mathbf{1}\cong{\it{\Gamma}}_{Y^{\prime}}\mathbf{1}\otimes L_{Z}\mathbf{1}$. ∎
We next prove some preliminary results regarding the localizing tensor ideals generated by Rickard idempotents. Recall that the localizing tensor ideal generated by a collection of objects $\mathbf{S}$ is denoted $\operatorname{Loc}^{\otimes}(\mathbf{S})$.
Lemma 5.8.
Let $\sigma$ be a comparative support with value in a Zariski space $X$ on a small monoidal triangulated category $\mathbf{K}^{c}$ satisfying (5.1). Let $V=\bigcup_{i\in I}V_{i}$. We have $\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{V_{i}}\mathbf{1}:i\in I)=\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{V}\mathbf{1})$.
Proof.
On one hand, since ${\it{\Gamma}}_{V_{i}}\mathbf{1}\otimes{\it{\Gamma}}_{V}\mathbf{1}\cong{\it{\Gamma}}_{V_{i}}\mathbf{1}$, it is clear that
$\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{V_{i}}\mathbf{1}:i\in I)\subseteq\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{V}\mathbf{1}).$
On the other hand, since ${\it{\Gamma}}_{V}$ is by definition in the localizing subcategory generated by $\Theta(V)$, it is enough to show that $\Theta(V)$ is contained in $\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{V_{i}}\mathbf{1}:i\in I).$ Suppose that $A\in\Theta(V)$. Since $\sigma(A)$ is closed and $X$ is a Zariski space, we can write $\sigma(X)$ as a finite union of irreducible closed sets, each of which has a unique generic point. In particular, there is some finite collection $V_{i_{1}},...,V_{i_{n}}$ such that
$\sigma(A)\subseteq V_{i_{1}}\cup...\cup V_{i_{n}}.$
Hence $A\in\Theta(V_{i_{1}}\cup...\cup V_{i_{n}})=\langle\Theta(V_{i_{1}}),...,\Theta(V_{i_{n}})\rangle$, using Lemma 5.5. But now note that
$\operatorname{Loc}(\langle\Theta(V_{i_{1}}),...,\Theta(V_{i_{n}})\rangle)\subseteq\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{V_{i_{1}}}\mathbf{1},...,{\it{\Gamma}}_{V_{i_{n}}}\mathbf{1}),$
since if $B\in\Theta(V_{i_{j}})$ then $B\otimes{\it{\Gamma}}_{V_{i_{j}}}\mathbf{1}\cong B$. The lemma now follows. ∎
Proposition 5.9.
Let $\sigma$ be a comparative support with value in a Zariski space $X$ on a small monoidal triangulated category $\mathbf{K}^{c}$ satisfying (5.1). Let $Y\subseteq X$ be some subset. Suppose $A\in\mathbf{K}$ satisfies ${\it{\Gamma}}_{x}A=0$ for all $x\not\in Y$. Then $A\in\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{y}\mathbf{1}\mid y\in Y)$.
Proof.
The proof closely follows [BHS23b, Theorem 3.22]. Set
$\mathbf{L}\coloneqq\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{y}\mathbf{1}\mid y\in Y).$
Define $Z\subseteq X$ to be
$Z\coloneqq\{x\in X\mid{\it{\Gamma}}_{\overline{\{x\}}}A\in\mathbf{L}\}.$
This is a specialization closed set. We claim that $Z=X$; this would prove the lemma, since then ${\it{\Gamma}}_{Z}\mathbf{1}\cong{\it{\Gamma}}_{X}\mathbf{1}\cong\mathbf{1}$, so
$A\cong{\it{\Gamma}}_{Z}A\in\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{\overline{\{x\}}}A\mid x\in Z)\subseteq\mathbf{L}.$
To prove the claim, suppose for the sake of contradiction that there exists $x\in X\setminus Z$; choose $x$ minimal with respect to specialization (which exists because $X$ is Noetherian). Then $S\coloneqq\overline{\{x\}}\setminus\{x\}\subseteq Z$ is specialization-closed (as is $\overline{\{x\}}$). Therefore we have
${\it{\Gamma}}_{S}\otimes{\it{\Gamma}}_{\overline{\{x\}}}\cong{\it{\Gamma}}_{S},$
and Lemma 5.7 implies that
${\it{\Gamma}}_{x}={\it{\Gamma}}_{\overline{\{x\}}}\otimes L_{S}.$
Applying the exact functor ${\it{\Gamma}}_{\overline{\{x\}}}$ to the distinguished triangle
${\it{\Gamma}}_{S}A\to A\to L_{S}A\to,$
we obtain
${\it{\Gamma}}_{S}A\to{\it{\Gamma}}_{\overline{\{x\}}}A\to{\it{\Gamma}}_{x}A\to.$
We claim that the first and third terms are in $\mathbf{L}$.
On one hand, by Lemma 5.8 we have
$\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{S}\mathbf{1})=\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{\overline{\{z\}}}\mathbf{1}:z\in S),$
which implies that ${\it{\Gamma}}_{S}A$ is in the localizing ideal generated by ${\it{\Gamma}}_{\overline{\{z\}}}A$ for all $z\in S$. Since $S\subseteq Z,$ each ${\it{\Gamma}}_{\overline{\{z\}}}A$ is in $\mathbf{L}$, so ${\it{\Gamma}}_{S}A\in\mathbf{L}$ as well. On the other hand, ${\it{\Gamma}}_{x}A\in\mathbf{L}$ because if $x\in Y$ then ${\it{\Gamma}}_{x}\mathbf{1}\in\mathbf{L}$ by definition, and if $x\not\in Y$ then ${\it{\Gamma}}_{x}A\cong 0$, which is also in $\mathbf{L}$.
Therefore ${\it{\Gamma}}_{\overline{\{x\}}}A\in\mathbf{L}$, implying that $x\in Z$, contradicting the assumption that $x\not\in Z$. Thus $Z\cong X$ and the claim is proved. ∎
Theorem 5.10.
Let $\sigma$ be a comparative support with value in a Zariski space $X$ on a small monoidal triangulated category $\mathbf{K}^{c}$ satisfying (5.1). Then the extension $\widetilde{\sigma}$ of $\sigma$ is faithful.
Proof.
The argument will be similar to the one given by Stevenson for the Balmer support in the commutative case [Ste13, Theorem 4.19(2)], although without the assumption that $\mathbf{K}$ arises as the homotopy category of a monoidal model category (this assumption was used in the analogue of Proposition 5.9, but was removed in [BHS23b, Theorem 3.22, Remark 3.24]).
First, note that Proposition 5.9 implies that
$\mathbf{K}=\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{x}\mathbf{1}:x\in X),$
since $\mathbf{1}$ is in $\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{x}\mathbf{1}:x\in X).$ From this, it follows that
$\operatorname{Loc}^{\otimes}(A)=\operatorname{Loc}^{\otimes}({\it{\Gamma}}_{x}A:x\in X).$
Therefore
$A\cong 0\iff\operatorname{Loc}^{\otimes}(A)=\{0\}\iff{\it{\Gamma}}_{x}A=0\;\;\forall\;\;x\in X\iff\widetilde{\sigma}(A)=\varnothing.$
In particular, since the Balmer support is always comparative, we may now deduce the following.
Corollary 5.11.
Let $\mathbf{K}^{c}$ be a small $\operatorname{Spc}$-Noetherian monoidal triangulated category. Then the extended Balmer support $\operatorname{Supp}^{\rm B}$ on $\mathbf{K}$ is faithful.
6. Faithfulness of extended tensorial support
Throughout this section, assume that $(X,\sigma)$ is a realizing tensorial support, and drop the assumption from the previous section that $\sigma$ is comparative. We will assume that $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian. In this section we determine when the extension $\widetilde{\sigma}$ is faithful, giving a characterization in terms of the universal map $\eta:X\to\operatorname{Spc}\mathbf{K}^{c}$.
Recall that by Proposition 2.6, since $\sigma$ is realizing, we have $\eta$ is injective, a fact that we use throughout the section. We note the following characterization of faithfulness of $\sigma$ in our setting in terms of the image of $\eta$; this lemma will not be used explicitly for the results of this section, but is helpful to keep in mind for the purpose of constructing examples.
Lemma 6.1.
Let $(X,\sigma)$ be a realizing tensorial support on a small $\operatorname{Spc}$-Noetherian monoidal triangulated category $\mathbf{K}^{c}$. Then $\sigma$ is faithful if and only if the image of $\eta$ contains all closed points of $\operatorname{Spc}\mathbf{K}^{c}$.
Proof.
Assume $\sigma$ faithful. If $\mathbf{P}$ is a closed point of $\operatorname{Spc}\mathbf{K}^{c}$, we can find an object $A$ such that $\operatorname{supp}^{\rm B}(A)=\{\mathbf{P}\}$ by Theorem 2.2 and the Noetherianity of $\operatorname{Spc}\mathbf{K}^{c}$. Since $\sigma(A)=\eta^{-1}(\operatorname{supp}^{\rm B}(A))$, and since $\sigma$ is faithful, there must be a point in the preimage of $\mathbf{P}$.
For the other direction, if all closed points are in the preimage of $\eta$, and if $A$ is in $\mathbf{K}^{c}$, then $\sigma(A)=\eta^{-1}(\operatorname{supp}^{\rm B}(A))$ is nonempty since $\operatorname{supp}^{\rm B}$ is faithful and closed sets of Noetherian topological spaces always contain closed points. ∎
We give examples below where, even though $\sigma$ is faithful, the extended support $\widetilde{\sigma}$ will fail to be faithful. Our goal this section will be to describe precisely when $\widetilde{\sigma}$ is faithful. In order to do that, we will introduce a convenient format for expressing the data of the underlying topological spaces of our supports.
Notation 6.2.
Let $X$ be a topological space. We will use the convention that the specialization poset of $X$ will be defined by $x\leq y$ if and only if $x\in\overline{\{y\}}$, for all $x$ and $y$ in $X$. We write the directed graph of the specialization poset by drawing an edge $x\to y$ if and only if $y\in\overline{\{x\}}$ and there does not exist another point $z\in X$ such that $\overline{\{x\}}\supsetneq\overline{\{z\}}\supsetneq\overline{\{y\}}$ (i.e., $y$ is a specialization of $x$ in the minimal sense).
Remark 6.3.
Recall that a space is called spectral if it is isomorphic to the spectrum of a commutative ring. Noetherian spectral spaces are completely characterized by their specialization posets: two Noetherian spectral spaces are homeomorphic if and only if their specialization posets are isomorphic [Ted17, Proposition 1.26]. Every finite poset arises as the specialization poset of a finite spectral space (see [Ted17, §3], which also surveys several methods to construct said rings). For any ring $R$ with Noetherian spectrum, we have a homeomorphism $\operatorname{Spec}R\cong\operatorname{Spc}\mathbf{K}^{c}$ where $\mathbf{K}=\operatorname{{\sf D}}(R)$ and $\mathbf{K}^{c}=\operatorname{{\sf D^{perf}}}(R)$ [Tho97] where $\operatorname{{\sf D}}(R)$ is the derived category of $R$-modules and $\operatorname{{\sf D^{perf}}}(R)$ is the perfect derived category of $R$-modules. Therefore any finite poset is the specialization poset of the Balmer spectrum of some small monoidal triangulated category, which we use to construct examples below.
We begin by giving an example of a tensorial realizing support whose extension is not faithful. Note that in the directed graphs of the specialization posets of $X$ and $\operatorname{Spc}\mathbf{K}^{c}$, closed points are always the end nodes and every connected component of the specialization poset contains at least one closed point.
Example 6.4.
Consider the following map of specialization posets:
${x}$${\widehat{y}}$${\widehat{z}}$${y}$${z}$
We can choose $\mathbf{K}^{c}$ such that the topological space associated to the right side, given by three points $\{x,y,z\}$ such that $y,z$ are closed points and $\overline{\{x\}}=\{x,y,z\}$, is homeomorphic to $\operatorname{Spc}\mathbf{K}^{c}$ by the discussion above. This spectral space is concretely realized as $\textnormal{Spec}\,R\cong\operatorname{Spc}\operatorname{{\sf D^{perf}}}(R)\cong\operatorname{Spc}\operatorname{{\sf D}}(R)^{c}$, where $R=\mathbb{Z}[\frac{1}{p}\mid p\neq 2,3]$. Now we consider an alternate support based on the topological space $X=\{\widehat{y},\widehat{z}\}$ consisting of two distinct closed points. We have the continuous map $\eta:X\to\operatorname{Spc}\mathbf{K}^{c}$ sending $\widehat{y}$ to $y$ and $\widehat{z}$ to $z$, as in the dotted arrows in the diagram above. This induces a support datum $\sigma$ based on $X$, where we define $\sigma(A)=\eta^{-1}(\operatorname{supp}^{\rm B}(A))$. This support is tensorial and realizing, which we can observe directly from the corresponding properties of $\operatorname{supp}^{\rm B}$. Consider the object
$A\coloneqq{\it{\Gamma}}_{x}\mathbf{1}={\it{\Gamma}}_{\mathcal{V}(x)}\mathbf{1}\otimes L_{\mathcal{Z}(x)}\mathbf{1}={\it{\Gamma}}_{\{x,y,z\}}\mathbf{1}\otimes L_{\{y,z\}}\mathbf{1}=L_{\{y,z\}}\mathbf{1}\in\mathbf{K}.$
This object has the property that $\operatorname{Supp}^{\rm B}(A)=\{x\}$, hence it is nonzero. We will see that $\widetilde{\sigma}(A)=\varnothing$. Let us examine the elements ${\it{\Gamma}}_{\widehat{y}}A$ and ${\it{\Gamma}}_{\widehat{z}}A$. We have
${\it{\Gamma}}_{\widehat{y}}\mathbf{1}$ $={\it{\Gamma}}_{\{\widehat{y}\}}\mathbf{1}\otimes L_{\{\widehat{z}\}}\mathbf{1},$
${\it{\Gamma}}_{\widehat{z}}\mathbf{1}$ $={\it{\Gamma}}_{\{\widehat{z}\}}\mathbf{1}\otimes L_{\{\widehat{y}\}}\mathbf{1}.$
Computing the ideals directly, we see by the definition of the support $\sigma$ that
$\Theta(\{\widehat{z}\})$ $=\Theta(\{z\}),$
$\Theta(\{\widehat{y}\})$ $=\Theta(\{y\}).$
This implies that
${\it{\Gamma}}_{\widehat{y}}\mathbf{1}\otimes A$ $={\it{\Gamma}}_{\{\widehat{y}\}}\mathbf{1}\otimes L_{\{\widehat{z}\}}\mathbf{1}\otimes L_{\{y,z\}}\mathbf{1}={\it{\Gamma}}_{\{y\}}\otimes L_{\{y,z\}}=0,$
${\it{\Gamma}}_{\widehat{z}}\mathbf{1}\otimes A$ $={\it{\Gamma}}_{\{\widehat{z}\}}\mathbf{1}\otimes L_{\{\widehat{y}\}}\mathbf{1}\otimes L_{\{y,z\}}\mathbf{1}={\it{\Gamma}}_{\{z\}}\otimes L_{\{y,z\}}=0,$
$\implies\widetilde{\sigma}(A)=\varnothing.$
This shows that we have a nontrivial support $(X,\sigma)$ which is faithful, exhibits realization and tensor product property, and yet its extension $\widetilde{\sigma}$ is not faithful.
The main result of this section characterizes exactly when the extension of a tensorial support is faithful; we deduce it from the following theorem, which tells us when the extended support of the ${\it{\Gamma}}_{\mathbf{P}}\mathbf{1}$ objects are empty.
Theorem 6.5.
Let $(X,\sigma)$ be a realizing faithful tensorial support on a small $\operatorname{Spc}$-Noetherian monoidal triangulated category $\mathbf{K}^{c}$ with extension $\widetilde{\sigma}$, and let $\eta:X\to\operatorname{Spc}\mathbf{K}^{c}$ be the universal map. Then for $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$, we have $\widetilde{\sigma}({\it{\Gamma}}_{\mathbf{P}}\mathbf{1})=\varnothing$ if and only if $\eta^{-1}(\overline{\{\mathbf{P}\}})$ does not have a unique generic point.
Recall that, in terms of the specialization poset, a generic point is a unique vertex which is the root of all points.
Proof.
Since $\eta$ is injective, if $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$ is in the image of $\eta$, it has a unique preimage; for ease of notation, denote this preimage by $\widehat{\mathbf{P}}$, if it exists.
First, $\widetilde{\sigma}({\it{\Gamma}}_{\mathbf{P}}\mathbf{1})\neq\varnothing$ if and only if there exists $\mathbf{Q}\in\operatorname{Spc}\mathbf{K}^{c}$ with preimage $\widehat{\mathbf{Q}}\in X$ such that ${\it{\Gamma}}_{\widehat{\mathbf{Q}}}\mathbf{1}\otimes{\it{\Gamma}}_{\mathbf{P}}\mathbf{1}\neq 0$. Since ${\it{\Gamma}}_{\widehat{\mathbf{Q}}}\mathbf{1}={\it{\Gamma}}_{\mathcal{V}(\widehat{\mathbf{Q}})}\mathbf{1}\otimes L_{\mathcal{Z}(\widehat{\mathbf{Q}})}\mathbf{1}$, by the tensor product property from Lemma 4.3, this is true if and only if $\mathbf{P}\in\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathcal{V}(\mathbf{Q})}\mathbf{1})$ and $\mathbf{P}\not\in\operatorname{Supp}^{\rm B}(L_{\mathcal{Z}(\{\widehat{\mathbf{Q}})\}}\mathbf{1})$. We will see that these two conditions force $\widehat{\mathbf{Q}}$ to be the unique generic point of $\eta^{-1}(\overline{\{\mathbf{P}\}})$.
By Lemma 4.2,
$\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathcal{V}(\widehat{\mathbf{Q}})}\mathbf{1})$ $=\Phi_{\operatorname{supp}^{\rm B}}(\Theta(\mathcal{V}(\widehat{\mathbf{Q}})))$
=$=\{\mathbf{R}\in\operatorname{Spc}\mathbf{K}^{c}:\exists A\in\Theta(\mathcal{V}(\widehat{\mathbf{Q}}))\text{ with }\mathbf{P}\in\operatorname{supp}^{\rm B}(A)\}$
=$=\{\mathbf{R}\in\operatorname{Spc}\mathbf{K}^{c}:\exists A\text{ with }\sigma(A)\subseteq\mathcal{V}(\widehat{\mathbf{Q}})\text{ and }\mathbf{P}\in\operatorname{supp}^{\rm B}(A)\}$
=$=\{\mathbf{R}\in\operatorname{Spc}\mathbf{K}^{c}:\exists A\text{ with }\eta^{-1}(\operatorname{supp}^{\rm B}(A))\subseteq\mathcal{V}(\widehat{\mathbf{Q}})\text{ and }\mathbf{P}\in\operatorname{supp}^{\rm B}(A)\}$
=$=\{\mathbf{R}\in\operatorname{Spc}\mathbf{K}^{c}:\eta^{-1}(\overline{\{\mathbf{P}\}})\subseteq\mathcal{V}(\widehat{\mathbf{Q}})\}.$
Hence $\mathbf{P}\in\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathcal{V}(\widehat{\mathbf{Q})}}\mathbf{1})$ if and only if $\eta^{-1}(\overline{\{\mathbf{P}\}})\subseteq\mathcal{V}(\widehat{\mathbf{Q}}).$ An analogous computation shows that $\mathbf{P}\not\in\operatorname{Supp}^{\rm B}(L_{\mathcal{Z}(\widehat{\mathbf{Q}})}\mathbf{1})$ if and only if $\eta^{-1}(\overline{\{\mathbf{P}\}})\not\subseteq\mathcal{Z}(\widehat{\mathbf{Q}}).$ But all points of $\mathcal{V}(\widehat{\mathbf{Q}})$ except for $\widehat{\mathbf{Q}}$ are also in $\mathcal{Z}(\widehat{\mathbf{Q}})$ (recall that we are assuming $X$ is $T_{0}$). To summarize, then, we have that $\mathbf{Q}\in\widetilde{\sigma}({\it{\Gamma}}_{\mathbf{P}}\mathbf{1})$ if and only if
$\widehat{\mathbf{Q}}\in\eta^{-1}(\overline{\{\mathbf{P}\}})\subseteq\mathcal{V}(\widehat{\mathbf{Q}}).$
But then we have
$\eta^{-1}(\overline{\{\mathbf{P}\}})=\mathcal{V}(\widehat{\mathbf{Q}}),$
that is, $\widehat{\mathbf{Q}}$ is a generic point of $\eta^{-1}(\overline{\{\mathbf{P}\}}).$
Corollary 6.6.
Let $(X,\sigma)$ be a realizing faithful tensorial support on a small $\operatorname{Spc}$-Noetherian monoidal triangulated category $\mathbf{K}^{c}$ with extension $\widetilde{\sigma}$. Then $\widetilde{\sigma}$ is faithful if and only if $\sigma$ is comparative.
Proof.
By Theorem 6.5, if $\widetilde{\sigma}$ is faithful then for each $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$, the closed subset $\eta^{-1}(\overline{\{\mathbf{P}\}})$ has a unique generic point. This is equivalent to $\sigma$ being comparative, by Proposition 5.3. For the other direction, if $\sigma$ is comparative, then we already know that $\widetilde{\sigma}$ is faithful, by Theorem 5.10. ∎
Example 6.7.
Consider the following example. We may choose $\mathbf{K}^{c}$ so that $\operatorname{Spc}\mathbf{K}^{c}$ is homeomorphic to the space of four points $\{a,b,c,d\}$ where $d$ is closed, $b$ and $c$ specialize to $d$, and $a$ is a generic point. Then we can consider $X$ to be the subspace consisting of preimages of $b$, $c$, and $d$; let $\eta:X\to\operatorname{Spc}\mathbf{K}^{c}$ be the inclusion map, so that $\eta$ induces the following map on the level of specialization posets:
${a}$${\widehat{b}}$${\widehat{c}}$${b}$${c}$${\widehat{d}}$${d}$
We have a realizing faithful tensorial support $\sigma$ induced by $\eta$. By Theorem 6.5, the extended support $\widetilde{\sigma}$ is not faithful, since the preimage $\eta^{-1}(\overline{\{a\}})=X$ does not have a unique generic point.
Example 6.8.
Consider Example 6.7 but where we now take $X$ to only have preimages of $b$ and $d$. Then $\eta$ induces the following map of specialization posets:
${a}$${\widehat{b}}$${b}$${c}$${\widehat{d}}$${d}$
In this case, $\widetilde{\sigma}$ is faithful by Theorem 6.5, since $\eta^{-1}(\overline{\{a\}})=X$, $\eta^{-1}(\overline{\{b\}})=\{\widehat{b},\widehat{d}\},$ $\eta^{-1}(\overline{\{c\}})=\{\widehat{d}\}$ and $\eta^{-1}(\overline{\{d\}})=\{\widehat{d}\}$ each have a unique generic point.
Remark 6.9.
In this situation, where $\sigma$ is faithful, tensorial, and realizing, there is a second possible extension of the support $\sigma$ other than $\widetilde{\sigma}$, namely,
$\widehat{\sigma}(A):=\eta^{-1}(\operatorname{Supp}^{\rm B}(A)).$
It is straightforward to verify that $\widehat{\sigma}$ is a tensorial extension of $\sigma$. However, $\widehat{\sigma}$ is faithful if and only if $\eta$ is a homeomorphism; this can be seen directly or deduced from Theorem 2.9. In fact, there is a containment
$\widehat{\sigma}(A)\subseteq\widetilde{\sigma}(A)$
for all $A$.
7. Faithfulness of extended support induced by a surjective map
In this section, we assume as usual that $\mathbf{K}$ is a big monoidal triangulated category. We also assume throughout the section that $\mathbf{K}^{c}$ is $\operatorname{Spc}$-Noetherian. We will prove that in this case support data on $\mathbf{K}^{c}$ induced by surjective maps admit faithful extensions.
We begin with a definition.
Definition 7.1.
Let $\rho:\operatorname{Spc}\mathbf{K}^{c}\to X$ be a surjective continuous map of topological spaces. We say that the map
$\sigma(A):=\overline{\rho(\operatorname{supp}^{\rm B}(A))}$
is the support induced by $\rho$.
Note that any comparative support is induced by a surjective map, since it is induced by the comparison map $\rho$ by Lemma 5.2, parts (1) and (2); the main difference between the assumptions in this section and those previously in Section 5 is that we have dropped the assumption that $\sigma$ is tensorial, but have added the assumption that $\mathbf{K}^{c}$ is $\operatorname{Spc}$-Noetherian. The purpose is for eventual application to the central cohomological support in the subsequent section (Section 8); for a finite tensor category $\mathbf{C}$, both the questions of Noetherianity of $\operatorname{Spc}\underline{\mathbf{C}}$ and the tensoriality of the central support are open.
Lemma 7.2.
Let $(X,\sigma)$ be the support induced by a surjective continuous map $\rho:\operatorname{Spc}\mathbf{K}^{c}\to X$ on a small $\operatorname{Spc}$-Noetherian monoidal triangulated category $\mathbf{K}^{c}$. Then $(X,\sigma)$ is a faithful and realizing support datum.
Proof.
It is straightforward that $(X,\sigma)$ is a support datum. Faithfulness follows from the fact that $\operatorname{supp}^{\rm B}$ is faithful (Example 2.4). Suppose $V\subseteq X$ is a closed subset. Then $\rho^{-1}(V)$ is a closed subset of $\operatorname{Spc}\mathbf{K}^{c}$, and so there exists $A\in\mathbf{K}^{c}$ with $\operatorname{supp}^{\rm B}(A)=\rho^{-1}(V)$. Surjectivity of $\rho$ then implies that $\sigma(A)=V$. ∎
Remark 7.3.
It is clear that a support induced by a surjective map need not be tensorial. For instance, if $\operatorname{Spc}\mathbf{K}^{c}$ is the topological space consisting of two distinct closed points, $X$ is the one-point space, and $\rho$ is the unique continuous map $\operatorname{Spc}\mathbf{K}^{c}\to X$. Then the support induced by $\rho$ is not tensorial: if $x$ is the unique point of $X$, and if $A$ and $C$ are two objects of $\mathbf{K}^{c}$ with disjoint nonempty supports (which exist, by Theorem 2.9), then we have $\sigma(A)=\{x\}=\sigma(C)$, but also have
$\bigcup_{B\in\mathbf{K}^{c}}\operatorname{supp}^{\rm B}(A\otimes B\otimes C)=\operatorname{supp}^{\rm B}(A)\cap\operatorname{supp}^{\rm B}(C)=\varnothing,$
hence $A\otimes B\otimes C\cong 0$ for all objects $B$. But then
$\bigcup_{B\in\mathbf{K}^{c}}\sigma(A\otimes B\otimes C)=\varnothing\subsetneq\sigma(A)\cap\sigma(B)=\{x\}.$
We now proceed in computing the extended Balmer supports of the Rickard idempotents associated to points of $X$.
Proposition 7.4.
Let $(X,\sigma)$ be the support induced by a surjective map $\rho$ on a small $\operatorname{Spc}$-Noetherian monoidal triangulated category $\mathbf{K}^{c}$, and let $V\subseteq X$ be a specialization-closed subset. We have
$\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{V}\mathbf{1})=\rho^{-1}(V).$
Proof.
Indeed, by Lemma 4.2,
$\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{V}\mathbf{1})$ $=\Phi_{\operatorname{supp}^{\rm B}}(\Theta(V))$
=$=\bigcup_{\{A\in\mathbf{K}^{c}:\sigma(A)\subseteq V\}}\operatorname{supp}^{\rm B}(A)$
=$=\bigcup_{\{A\in\mathbf{K}^{c}:\overline{\rho(\operatorname{supp}^{\rm B}(A))}\subseteq V\}}\operatorname{supp}^{\rm B}(A)$
=$=\bigcup_{\{A\in\mathbf{K}^{c}:\rho(\operatorname{supp}^{\rm B}(A))\subseteq V\}}\operatorname{supp}^{\rm B}(A)$
=$=\bigcup_{\{A\in\mathbf{K}^{c}:\operatorname{supp}^{\rm B}(A)\subseteq\rho^{-1}(V)\}}\operatorname{supp}^{\rm B}(A)$
$=\rho^{-1}(V).$
The last line follows since Balmer support satisfies realization, using the Noetherianity assumption.
Corollary 7.5.
Let $(X,\sigma)$ be the support induced by a surjective map $\rho$ on a small $\operatorname{Spc}$-Noetherian monoidal triangulated category $\mathbf{K}^{c}$, and let $x\in X$. We have
$\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{x}\mathbf{1})=\rho^{-1}(x).$
Proof.
Let $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$. Since ${\it{\Gamma}}_{x}\mathbf{1}={\it{\Gamma}}_{\mathcal{V}(x)}\mathbf{1}\otimes L_{\mathcal{Z}(x)}\mathbf{1}$, and since the extended Balmer support is tensorial by Proposition 4.4, it is clear that if $\mathbf{P}$ is in $\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{x})$ then $\mathbf{P}\in\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathcal{V}(x)})$ and $\mathbf{P}\not\in\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{\mathcal{Z}(x)}\mathbf{1})$ (recall Lemma 4.2). Since $\mathcal{V}(x)\cap(X\backslash\mathcal{Z}(x))=\{x\},$ the proposition follows from Proposition 7.4. ∎
Since $\sigma$ is not tensorial, we cannot assume that $\widetilde{\sigma}$ is actually an extension of $\sigma$, at this point, since Proposition 3.8 requires tensoriality. However, we are able to prove the following description of $\widetilde{\sigma}$, which will then allow us to describe precisely when $\widetilde{\sigma}$ agrees with $\sigma$ on compact objects (i.e., is an extension).
Theorem 7.6.
Let $(X,\sigma)$ be the support datum induced by a surjective map $\rho$ on a $\operatorname{Spc}$-Noetherian small monoidal triangulated category $\mathbf{K}^{c}$. Then
$\rho(\operatorname{Supp}^{\rm B}(A))=\widetilde{\sigma}(A)$
for all $A\in\mathbf{K}$. As a consequence, $\widetilde{\sigma}$ is faithful. It is an extended support if and only if the map $\rho$ is closed.
Proof.
For the containment $\subseteq$, let $x=\rho(\mathbf{P})$ for some $\mathbf{P}\in\operatorname{Supp}^{\rm B}(A).$ Then $\mathbf{P}\in\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{x}\mathbf{1})$ by Corollary 7.5. Since $\operatorname{Supp}^{\rm B}$ is tensorial, this implies $\mathbf{P}\in\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{x}\otimes A)$. This implies that ${\it{\Gamma}}_{x}\otimes A\not=0$, and hence $x\in\widetilde{\sigma}(A)$.
For the containment $\supseteq$, if $x\in\widetilde{\sigma}(A)$, then we have ${\it{\Gamma}}_{x}\otimes A\not=0$. Since $\operatorname{Supp}^{\rm B}$ is faithful by Corollary 5.11, $\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{x}\otimes A)=\operatorname{Supp}^{\rm B}({\it{\Gamma}}_{x})\cap\operatorname{Supp}^{\rm B}(A)=\rho^{-1}(x)\cap\operatorname{Supp}^{\rm B}(A)$ is nonempty. In particular, there exists $\mathbf{P}$ in this intersection. But then $x=\rho(\mathbf{P})$, and this $\mathbf{P}$ is in $\operatorname{Supp}^{\rm B}(A)$, so $x\in\rho(\operatorname{Supp}^{\rm B}(A))$.
If $A$ is compact, we have now shown that
$\widetilde{\sigma}(A)$ $=\rho(\operatorname{Supp}^{\rm B}(A))$
$=\rho(\operatorname{supp}^{\rm B}(A)).$
Clearly, faithfulness of $\widetilde{\sigma}$ follows then from the faithfulness of $\operatorname{Supp}^{\rm B}$.
If $\rho$ is closed, we have
$\widetilde{\sigma}(A)=\rho(\operatorname{Supp}^{\rm B}(A))=\rho(\operatorname{supp}^{\rm B}(A))=\sigma(A)$
for all $A$ compact, so $\widetilde{\sigma}$ extends $\sigma$. On the other hand, if $\rho$ is not closed, then there exists some closed set $\operatorname{supp}^{\rm B}(A)$ with
$\widetilde{\sigma}(A)=\rho(\operatorname{Supp}^{\rm B}(A))=\rho(\operatorname{supp}^{\rm B}(A))\subsetneq\overline{\rho(\operatorname{supp}^{\rm B}(A))}=\sigma(A)$
for some compact $A$, using the fact that $\operatorname{Supp}^{\rm B}$ extends $\operatorname{supp}^{\rm B}$ and that $\operatorname{supp}^{\rm B}$ satisfies realization, since $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian (Example 2.4). ∎
8. Applications to stable categories of finite tensor categories
In this section, let $\mathbf{C}$ be a finite tensor category as in [EGNO15]. That is, $\mathbf{C}$ is an abelian $𝕜\Bbbk$-linear rigid monoidal category for some algebraically-closed field $𝕜\Bbbk$, each space of morphisms between objects of $\mathbf{C}$ is finite-dimensional, $\mathbf{C}$ has finitely many simple objects, all objects of $\mathbf{C}$ have finite length, and the unit object $\mathbf{1}$ is simple. In this case, the stable category $\underline{\mathbf{C}}$ of $\mathbf{C}$ is defined as having the same objects as $\mathbf{C}$, and morphism spaces
$\operatorname{Hom}_{\underline{\mathbf{C}}}(A,B):=\operatorname{Hom}_{\mathbf{C}}(A,B)/\operatorname{PHom}_{\mathbf{C}}(A,B),$
where $\operatorname{PHom}_{\mathbf{C}}(A,B)$ denotes the collection of morphisms $A\to B$ in $\mathbf{C}$ which factor through a projective object. The stable category $\underline{\mathbf{C}}$ is monoidal triangulated [Hap88]. In fact, $\underline{\mathbf{C}}$ is a small monoidal triangulated category: it appears as the compact part of the rigidly-compactly-generated monoidal triangulated category $\underline{\operatorname{Ind}(\mathbf{C})}$ [NVY24, Appendix A]. We will denote
$\mathbf{K}:=\underline{\operatorname{Ind}(\mathbf{C})},\;\;\mathbf{K}^{c}:=\underline{\mathbf{C}}.$
Since $\mathbf{C}$ has finitely many simple objects, $\underline{\mathbf{C}}$ is finitely-generated: the generator is the image of the direct sum of the simples.
Example 8.1.
One of the motivating examples of finite tensor categories is the category of finite-dimensional representations of a finite-dimensional Hopf algebra $H$. In this case, the category $\operatorname{Ind}(\mathbf{C})$ is the category of all representations of $H$.
Recall that the cohomology ring $R^{\bullet}$ of $\mathbf{C}$ is the $𝕜\Bbbk$-algebra
$R^{\bullet}:=\bigoplus_{i\geq 0}\operatorname{Ext}^{i}_{\mathbf{C}}(\mathbf{1},\mathbf{1})$
under the Yoneda product. The cohomology can also be described via the triangulated structure of $\mathbf{K}^{c}$ (see e.g. [CTVEZ03, Proposition 2.6.2]):
$R^{\bullet}\cong\bigoplus_{i\geq 0}\operatorname{Hom}_{\mathbf{K}^{c}}(\mathbf{1},\Sigma^{i}\mathbf{1}).$
The categorical center $C^{\bullet}\subseteq R^{\bullet}$ was introduced in [NVY24], and is defined as the subalgebra of $R^{\bullet}$ generated by homogeneous elements $g:\mathbf{1}\to\Sigma^{i}\mathbf{1}$ such that the diagrams
[…]
commute for all simple objects $A$.
By composing shifts of morphisms as in [NVY24, Section 1.3], there is a canonical action of the categorical center $C^{\bullet}$ on each algebra $\operatorname{Hom}^{\bullet}(A,A):=\bigoplus_{i\geq 0}\operatorname{Hom}_{\mathbf{K}^{c}}(A,\Sigma^{i}A).$ We say that $\mathbf{K}^{c}$ satisfies the weak finite generation condition, or (wfg) condition, if each $\operatorname{Hom}^{\bullet}(A,A)$ is finitely-generated as a $C^{\bullet}$ module. It would be a consequence of the Etingof–Ostrik conjecture (see [EO04]) that $\mathbf{K}^{c}=\underline{\mathbf{C}}$ would satisfy the (wfg) condition, for each finite tensor category $\mathbf{C}$ (see [NVY24, Section 1.5]).
We have the central cohomological support defined by
$\operatorname{supp}^{\rm C}(A):=\{\mathfrak{p}\in\operatorname{Proj}C^{\bullet}:I(A)\subseteq\mathfrak{p}\}$
where $I(A)$ is the annihilator ideal of $\operatorname{Hom}^{\bullet}(A,A)$ in $C^{\bullet}$. The central cohomological support is a support datum (see [NVY24, Proposition 6.1.1]). As long as $\mathbf{K}^{c}$ satisfies the (wfg) condition, the central cohomological support is faithful, see the discussion in [NVY24, Corollary 7.1.3], which uses the result [BPW21, Corollary 4.2] that connects dimension of a support variety with the complexity of the module.
When $\mathbf{K}^{c}$ satisfies the (wfg) condition, there exists a surjective continuous map $\operatorname{Spc}\mathbf{K}^{c}\to\operatorname{Proj}C^{\bullet}$ given by
$\rho(\mathbf{P}):=\langle g\in C^{\bullet}:g\text{ homogeneous,}\operatorname{cone}(g)\not\in\mathbf{P}\rangle$
for any prime ideal $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$, see [NVY24, Corollary 7.1.3, Theorem 7.2.1], a noncommutative analogue of Balmer’s comparison map [Bal10]. If the central support $\operatorname{supp}^{\rm C}$ is tensorial, then we have the universal map
[…]
In this case $\rho\circ\eta=\operatorname{id}_{\operatorname{Proj}C^{\bullet}}$ and $(\eta\circ\rho)(\mathbf{P})\subseteq\mathbf{P}$ for all $\mathbf{P}\in\operatorname{Spc}\mathbf{K}^{c}$ by [NVY24, Proposition 8.1.1]. Whether or not $\operatorname{supp}^{\rm C}$ is tensorial in general is still an open question.
Notation 8.2.
The extension of the central cohomological support $\operatorname{supp}^{\rm C}$ will be denoted $\operatorname{Supp}^{\rm C}$.
Before stating our main theorem giving conditions under which the central cohomological support satisfies the necessary assumptions given in the previous sections for there to exist a faithful extension, we prove some preliminary results.
Proposition 8.3.
Let $\mathbf{K}^{c}=\underline{\mathbf{C}}$ be the stable category of a finite tensor category $\mathbf{C}$ satisfying weak finite generation, and suppose that the categorical center $C^{\bullet}$ is finitely-generated. Let $A\in\mathbf{K}$ and let $G$ be the image of the direct sum of the simple objects of $\mathbf{C}$ in $\mathbf{K}^{c}.$ Then $\operatorname{supp}^{\rm C}(A\otimes G\otimes A)=\operatorname{supp}^{\rm C}(A).$
Proof.
We have the containment
$\operatorname{supp}^{\rm C}(A\otimes G\otimes A)\subseteq\operatorname{supp}^{\rm C}(A)$
by [NVY24, Proposition 6.1.1(e)]. Suppose that this containment is strict. Then there exists some homogeneous maximal ideal $\mathfrak{m}\in\operatorname{supp}^{\rm C}(A)\subseteq\operatorname{Proj}C^{\bullet}$ which is not in $\operatorname{supp}^{\rm C}(A\otimes G\otimes A)$ since in schemes of locally finite type over a field $𝕜\Bbbk$, closed subsets are determined by their closed points (see e.g. [GW10, Proposition 3.35]). Take a collection of generators $g_{1},...,g_{n}$ of $\mathfrak{m}$. Set $B=\operatorname{cone}(g_{1})\otimes...\otimes\operatorname{cone}(g_{n})$. We have $\operatorname{supp}^{\rm C}(B)=\{\mathfrak{m}\}=\operatorname{supp}^{\rm C}(A\otimes B)$ by [NVY24, Proposition 7.1.1(a)].
Now note that for any objects $D$ and $E$ in $\mathbf{K}^{c}$, we have a distinguished triangle
$D\otimes\mathbf{1}\otimes E\xrightarrow{\operatorname{id}_{D}\otimes g_{i}\otimes\operatorname{id}_{E}}\mathbf{1}\to D\otimes\operatorname{cone}(g_{i})\otimes E\to\Sigma(D\otimes\mathbf{1}\otimes E).$
It follows that $\operatorname{supp}^{\rm C}(D\otimes\operatorname{cone}(g_{i})\otimes E)\subseteq\operatorname{supp}^{\rm C}(D\otimes E)$. Since $B$ is a tensor product of all of the $\operatorname{cone}(g_{i})$’s, we also have $\operatorname{supp}^{\rm C}(D\otimes B\otimes E)\subseteq\operatorname{supp}^{\rm C}(D\otimes E),$ for any objects $D$ and $E$.
By the above discussion and [NVY24, Proposition 7.1.1(d)], we have
$W_{C}((A\otimes B)\otimes G\otimes(A\otimes B))$ $\subseteq W_{C}(A\otimes G\otimes A\otimes B)$
$=W_{C}(A\otimes G\otimes A)\cap W_{C}(B)$
==$\varnothing$.= ∅ .
It follows that $(A\otimes B)\otimes G\otimes(A\otimes B)\cong 0$ by the faithfulness of the central cohomological support, and hence $A\otimes B\cong 0$ by [NVY22b, Proposition 4.1.1]. But this is a contradiction, since $A\otimes B$ has a nontrivial support variety. ∎
Proposition 8.4.
Let $\mathbf{K}^{c}=\underline{\mathbf{C}}$ be the stable category of a finite tensor category $\mathbf{C}$ satisfying weak finite generation, and suppose that the categorical center $C^{\bullet}$ is finitely-generated. We have $\rho(\operatorname{supp}^{\rm B}(A))=\operatorname{supp}^{\rm C}(A)$ for all $A\in\mathbf{K}^{c}$.
Proof.
We know by [NVY24, Corollary 6.2.5(b)] that $\rho(\operatorname{supp}^{\rm B}(A))\subseteq\operatorname{supp}^{\rm C}(A)$.
For the other direction, suppose $\mathfrak{p}\in\operatorname{supp}^{\rm C}(A)$. We first show that $\mathfrak{p}\in\overline{\rho(\operatorname{supp}^{\rm B}(A))}.$ To do this, we must produce a prime $\mathbf{P}\in\operatorname{supp}^{\rm B}(A)$ such that $\mathfrak{p}\supseteq\rho(\mathbf{P})$. Consider the thick ideal
$\mathbf{I}:=\langle\operatorname{cone}(g):g\not\in\mathfrak{p}\rangle,$
and the collection of objects
$\mathbf{M}:=\{A\otimes G\otimes A...\otimes G\otimes A\},$
where $G$ is the sum of simple objects. So long as $\mathbf{I}$ and $\mathbf{M}$ are disjoint, we may produce a prime ideal containing $\mathbf{I}$ which intersects $\mathbf{M}$ trivially, by [NVY23, Lemma A.1.1], since $\mathbf{M}$ has the property that given any two objects $B$ and $D$, there exists an object $C$ such that $B\otimes C\otimes D\in\mathbf{M}$ (of course, we have constructed it so that $D$ can be chosen as $G$). Indeed, $\mathbf{I}$ and $\mathbf{M}$ must be disjoint, since by [NVY24, Proposition 7.1.1(d)] we have $\operatorname{supp}^{\rm C}(\operatorname{cone}(g))=Z(g)$ for any $g\in C^{\bullet}$, hence
$\operatorname{supp}^{\rm C}(B)\subseteq Z(\{g:g\not\in\mathfrak{p}\})$
for any $B\in\mathbf{I}$. Therefore, $\mathfrak{p}\not\in\operatorname{supp}^{\rm C}(B)$ for any $B\in\mathbf{I}$. Since $\operatorname{supp}^{\rm C}(A\otimes G\otimes A\otimes...\otimes G\otimes A)=\operatorname{supp}^{\rm C}(A)$ by Proposition 8.3, and $\mathfrak{p}\in\operatorname{supp}^{\rm C}(A)$, it follows that $\mathbf{I}\cap\mathbf{M}$ is empty. Thus, by [NVY23, Lemma A.1.1], we can produce a prime ideal $\mathbf{P}$ containing $\mathbf{I}$ and disjoint from $\mathbf{M}$. Since $\mathbf{P}$ contains $\mathbf{I}$, we have
$\rho(\mathbf{P})$ $=\langle g\in C^{\bullet}:g\text{ homogeneous},\operatorname{cone}(g)\not\in\mathbf{P}\rangle$
$\subseteq\langle g\in C^{\bullet}:g\text{ homogeneous },g\in\mathfrak{p}\rangle$
$=\mathfrak{p}.$
That is, $\mathfrak{p}$ is in the closure of $\rho(\mathbf{P})$. But now note that since $\mathbf{P}$ is disjoint from $\mathbf{M}$, it in particular does not contain $A$, that is, $\mathbf{P}\in\operatorname{supp}^{\rm B}(A)$.
Now, since $\rho$ is surjective, there exists $\mathbf{Q}$ with $\rho(\mathbf{Q})=\mathfrak{p}$. By the proof of [NVY24, Theorem 7.2.1], such a $\mathbf{Q}$ necessarily is disjoint from $\mathbf{I}$ and $\mathbf{M}$ as given above, and hence we can indeed choose $\mathbf{P}$ above which contains $\mathbf{Q}$. Since $\mathbf{Q}\subseteq\mathbf{P}$, we have $\mathbf{Q}\in\operatorname{supp}^{\rm B}(A)$, and so $\mathfrak{p}\in\rho(\operatorname{supp}^{\rm B}(A)).$
Immediately from Proposition 8.4, we deduce:
Corollary 8.5.
Let $\mathbf{K}^{c}=\underline{\mathbf{C}}$ be the stable category of a finite tensor category $\mathbf{C}$ satisfying weak finite generation, and suppose that the categorical center $C^{\bullet}$ is finitely-generated and that $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian. Then the map $\rho$ is a closed map.
We have now completed the work necessary to state our main results for central support on the stable category of a finite tensor category.
Theorem 8.6.
Let $\mathbf{K}^{c}=\underline{\mathbf{C}}$ be the stable category of a finite tensor category $\mathbf{C}$ satisfying the weak finite-generation condition. Then if $\mathbf{K}^{c}$ satisfies either of the following two conditions, the central support $\operatorname{supp}^{\rm C}$ admits a faithful extension:
  1. (1)
    $\operatorname{Proj}C^{\bullet}$ is Noetherian and $\operatorname{supp}^{\rm C}$ is tensorial;
  2. (2)
    $C^{\bullet}$ is finitely-generated and $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian.
Proof.
(1): If $\operatorname{supp}^{\rm C}$ is tensorial and $\operatorname{Proj}C^{\bullet}$ Noetherian, then the map $\rho$ defined above satisfies the axioms of a comparison map given in Definition 5.1, so that $\operatorname{supp}^{\rm C}$ is a comparative support and we can deduce that it has a faithful extension, by Theorem 5.10.
(2): We know that central support $\operatorname{supp}^{\rm C}$ is induced by the surjective map $\rho$ by Proposition 8.4. Hence if $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian, the result follows from Theorem 7.6. ∎
Remark 8.7.
Benson–Iyengar–Krause constructed extensions in [BIK08] for supports arising from an arbitrary graded-commutative Noetherian ring $R$ acting on a triangulated category. However, this support is based on the homogeneous prime spectrum of $R$ rather than $\operatorname{Proj}R$, in other words, the irrelevant ideal (if the ring is $\mathbb{N}$-graded) is allowed as a point. Other than this difference, the central cohomological support fits into their framework, and our extension is defined in the same way as theirs (indeed, their extension serves as one of the primary models for the extension in this paper). In [BIK08, Theorem 5.2], they prove that their extension is faithful. This does not immediately imply that the extension for the $\operatorname{Proj}$-based version is faithful, however, since a priori their theorem still allows for a nonzero object in $\mathbf{K}$ to have support equal to the irrelevant ideal, which would translate to empty $\operatorname{Supp}^{\rm C}$-support. We may therefore view Theorem 8.6 as a strengthening of the faithfulness result of Benson–Iyengar–Krause under tensoriality or $\operatorname{Spc}$-Noetherianity assumptions.
Example 8.8.
[…]
Pevtsova and Witherspoon prove in [PW15, Theorem 1.1] that the cohomological support of $\mathbf{K}^{c}$ is tensorial, and hence we have a continuous map
[…]
$𝕜\rho:\operatorname{Spc}\mathbf{K}^{c}\to\operatorname{Proj}\operatorname{H}^{\bullet}(A,\Bbbk)$
recalled above is a left inverse to $\eta$, and is closed by Proposition 8.4. By the classification of thick ideals for $\mathbf{K}^{c}$, also due to Pevtsova–Witherspoon [PW15, Theorem 1.2], $\rho$ must be injective, and hence $\eta$ and $\rho$ are inverse homeomorphisms. Indeed, using the explict formula for $\rho$, if $\rho(\mathbf{P})=\rho(\mathbf{Q})$ for $\mathbf{P},\mathbf{Q}\in\operatorname{Spc}\mathbf{K}^{c}$, then it would follow that for all $𝕜g\in\operatorname{H}^{\bullet}(A,\Bbbk)$, we would have $\operatorname{cone}(g)\in\mathbf{P}\Leftrightarrow\operatorname{cone}(g)\in\mathbf{Q}$, which would imply that the supports of $\mathbf{P}$ and $\mathbf{Q}$ would be the same, by [NVY24, Proposition 7.1.1], a contradiction. Therefore
[…]
These results allow us to apply Theorem 8.6, since $\operatorname{Proj}C^{\bullet}$ is Noetherian, $\operatorname{supp}^{\rm C}$ is tensorial, $C^{\bullet}$ is finitely-generated, and $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian. We obtain that $\operatorname{supp}^{\rm C}$ admits a faithful extension to all of $\mathbf{K}$. Furthermore, since $\operatorname{supp}^{\rm C}$ identifies with the Balmer support, by Proposition 4.4 we have tensoriality of $\operatorname{Supp}^{\rm C}$: the question posed by Pevtsova–Witherspoon, asking whether there exists an extension of $\operatorname{supp}^{\rm C}$ satisfying a tensor product property, has a positive answer.
We note that this problem was first solved in [NP23a] using an alternate method. Negron–Pevtsova construct a different extended support, termed hypersurface support, and prove in [NP23a, Theorem 7.9] that this hypersurface support is a faithful extended support satisfying a version of the tensor product property. However, for hypersurface support to be defined, one needs to have a smoothly integrable Hopf algebra. We emphasize that the techniques we use to recover this result above make no reference to the smooth integration, and that by Proposition 4.4 and Corollary 5.11 we expect existence of a faithful and tensorial extended support to be a general phenomena in monoidal triangular geometry, and not specific to smoothly integrable Hopf algebras.